### v2

# Constrained Denoising, Empirical Bayes, and Optimal Transport

Abstract
In the statistical problem of denoising, Bayes and empirical Bayes methods can “overshrink” their output relative to the latent variables of interest. This work is focused on constrained denoising problems which mitigate such phenomena. At the oracle level, i.e., when the latent variable distribution is assumed known, we apply tools from the theory of optimal transport to characterize the solution to (i) variance-constrained, (ii) distribution-constrained, and (iii) general-constrained denoising problems. At the empirical level, i.e., when the latent variable distribution is not known, we use empirical Bayes methodology to estimate these oracle denoisers. Our approach is modular, and transforms any suitable (unconstrained) empirical Bayes denoiser into a constrained empirical Bayes denoiser. We prove explicit rates of convergence for our proposed methodologies, which both extend and sharpen existing asymptotic results that have previously considered only variance constraints. We apply our methodology in two applications: one in astronomy concerning the relative chemical abundances in a large catalog of red-clump stars, and one in baseball concerning minor- and major league batting skill for rookie players.
Keywords: constrained Bayes estimation; $G$-modeling; Gaussian mixture model; errors-in-variables regression; latent variable model; (smoothed) nonparametric maximum likelihood estimation; Wasserstein space
1 Introduction
1.1 Problem Statement
This work is focused on the statistical problem of denoising, which concerns the following model for the joint distribution for a pair of random variables $(\Theta,Z)\in\mathbb{R}^{m}\times\mathbb{R}^{d}$ for $m\geq 1$ and $d\geq 1$:
$\Theta\sim G,\qquad\mbox{and}\qquad(Z\mid\Theta=\theta)\sim P_{\theta}.$ (1.1)
Here, $G$ is an unknown distribution in $\mathbb{R}^{m}$ and $\{P_{\theta}\}_{\theta}$ is a known family of probability distributions on $\mathbb{R}^{d}$. We assume that there exist $(\Theta_{1},Z_{1}),\ldots,(\Theta_{n},Z_{n})$ which are independent identically distributed (i.i.d.) pairs from the above model; we refer to $\Theta_{1},\ldots,\Theta_{n}$ as the latent variables, we refer to $Z_{1},\ldots,Z_{n}$ as the observations, and denoising is the task of estimating/predicting the latent variables from the observations.
More precisely, the goal of denoising is to construct a function $\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}$ such that the risk $\mathbb{E}[\|\delta(Z)-\Theta\|^{2}]$ is small, with expectation taken over the joint distribution of $(\Theta,Z)$. If the distribution $G$ is known, then the best choice is the posterior mean $\delta_{\mathcal{B}}(z):=\mathbb{E}[\Theta\,|\,Z=z]$, called the Bayes denoiser (with respect to the squared error loss). If the distribution $G$ is not known, then it is often still possible to construct some function $\hat{\delta}_{\mathcal{B}}$, called an empirical Bayes (EB) denoiser (Robbins, 1956; Efron, 2019), which achieves a risk that is nearly equal to that of the Bayes denoiser.
This capability suggests a strategy of undeniable appeal for practical data analysis on the latent variables $\Theta_{1},\ldots,\Theta_{n}$: First design the analysis as if the latent variables $\Theta_{1},\dotsc,\Theta_{n}$ were available, then apply this same analysis to the denoised observations $\hat{\delta}_{\mathcal{B}}(Z_{1}),\ldots,\hat{\delta}_{\mathcal{B}}(Z_{n})$. The reader may notice, however, that this plug-in approach can be problematic; the distribution of $\hat{\delta}_{\mathcal{B}}(Z_{1}),\ldots,\hat{\delta}_{\mathcal{B}}(Z_{n})$ need not be close at all to the distribution of $\Theta_{1},\ldots,\Theta_{n}$, so downstream analyses based on this strategy need not be even consistent. This was summarily described by Louis (1984) (where his “parameters” are our “latent variables”):
“In standard Bayes and empirical Bayes component decision problems, estimating individual parameters is the primary goal. In […] comparisons of histograms of estimates, however, the primary goal is to produce parameter estimates that can be considered as an ensemble. For example, the histogram of estimates should be a good estimate of the histogram of parameters.”
Similar considerations have been noted by many authors over the last hundred years, e.g., Eddington (1940); Louis (1984); Ghosh (1992); Loredo (2007); Bloom et al. (2017); Agarwal et al. (2020); Chen et al. (2025), and in textbooks [Raudenbush and Bryk, 2002, p. 88; Rao and Molina, 2015, p. 316].
Figure 1.1: Denoising in a simulated data set. The latent variables come from a two-component Gaussian mixture model, and the observations come from a Gaussian likelihood. We show the latent variables and the observations (first column), the Bayes and EB denoisers (second column), the variance-constrained Bayes and EB denoisers (third column), and the distribution-constrained Bayes and EB denoisers (fourth column). In each plot, we also show the contour for two standard deviations around each component (blue) for reference.
The fundamental problem with the plug-in approach is that $\hat{\delta}_{\mathcal{B}}$ achieving nearly the optimal risk does not imply that the distribution of $\hat{\delta}_{\mathcal{B}}(Z_{1}),\ldots,\hat{\delta}_{\mathcal{B}}(Z_{n})$ must be close to the distribution of $\Theta_{1},\ldots,\Theta_{n}$. Instead, it often implies that the former is “overshrunk” compared to the latter. This can be seen already at the population level, since $\hat{\delta}_{\mathcal{B}}$ targets $\delta_{\mathcal{B}}$, and the law of total covariance yields the following, where $\prec$ refers to the positive semi-definite order:
$\textnormal{Cov}(\delta_{\mathcal{B}}(Z))=\textnormal{Cov}(\Theta)-\mathbb{E}[\textnormal{Cov}(\Theta\,|\,Z)]\preceq\textnormal{Cov}(\Theta).$ (1.2)
As a concrete example of this phenomenon, consider the simple setting where $\Theta\sim\mathcal{N}(0,1)$ and $P_{\theta}=\mathcal{N}(\theta,1)$, and note that we have $\delta_{\mathcal{B}}(Z)\equiv Z/2\sim\mathcal{N}(0,\nicefrac{{1}}{{2}})\neq\mathcal{N}(0,1)$. As another example, consider the simulation given in the first two columns of Figure 1.1 where the latent variables come from a two-component Gaussian mixture model (with component variance $\tau^{2}I_{2}$ with $\tau^{2}=0.1$) and the observations come from a Gaussian likelihood (with likelihood variance $\sigma^{2}I_{2}$ with $\sigma^{2}=1.0$). From $n=2{,}000$ such observations, we observe that the Bayes and EB denoisers both shrink the observations too aggressively. (Presently, we compute the EB denoiser using the $G$-modeling approach of the smooth NPMLE from Magder and Zeger (1996) as described in Example 4.2 with variance lower bound $\tau^{2}I_{2}$.)
The goal of this work is to develop methodology for constructing fully-data-driven denoisers $\hat{\delta}$ such that the distribution of $\hat{\delta}(Z_{1}),\ldots,\hat{\delta}(Z_{n})$ must be, in some sense, close to the distribution of $\Theta_{1},\ldots,\Theta_{n}$, and which achieve a risk that is nearly equal to that of the optimal constrained denoiser. The other panels of Figure 1.1 provide a preview of our methodology, which we will precisely define later.
1.2 Classical Denoising and Three Constrained Denoising Problems
We now describe the classical denoising problems and three constrained variants that are of interest in this work. First, we consider the (unconstrained) Bayes denoising problem,
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right],$ ($\mathcal{B}$)
whose solution is the posterior mean $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$, and we next add suitable constraints.
Variance-Constrained Denoising. The following problem has been studied by Louis (1984); Ghosh (1992), and several others, and is precisely given as
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\qquad\textnormal{s.t.}\qquad\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta]\quad\textnormal{and}\quad\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta).$ ($\mathcal{VCB}$)
Problem ($\mathcal{VCB}$) aims to find a denoiser $\delta$ that minimizes the Bayes risk subject to the constraint that the first two moments of $\delta$ match those of the latent variable $\Theta$, thus mitigating the shrinkage of (1.2).
Distribution-Constrained Denoising. The second problem of interest has been recently studied by García-Trillos and Sen (2024), and it is given as
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\qquad\textnormal{s.t.}\qquad\delta(Z)\overset{\mathcal{D}}{=}\Theta.$ ($\mathcal{DCB}$)
Here, rather than enforcing constraints on the first two moments, we require that the distribution of $\delta(Z)$ matches exactly the distribution of the latent variable $\Theta$.
General-Constrained Denoising. We may accommodate general constraints via the problem
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\qquad\textnormal{s.t.}\qquad\mathbb{E}[\psi_{\ell}(\delta(Z))]=\mathbb{E}[\psi_{\ell}(\Theta)]\quad\textnormal{ for all }1\leq\ell\leq k$ ($\mathcal{GCB}$)
where $\psi_{1},\ldots,\psi_{k}:\mathbb{R}^{m}\to\mathbb{R}$ is some prescribed set of measurable functions. This formulation is general enough to cover a wide variety of problems; for example, it encompasses ($\mathcal{VCB}$) when $m=1$ by taking $\psi_{1}(z):=z$ and $\psi_{2}(z):=z^{2}$, it also encompasses ($\mathcal{DCB}$) in a limiting sense when $\psi_{1},\psi_{2},\ldots$ are taken to be a suitable sequence of bounded measurable functions. The added generality is useful in some novel applications and substantially extends the scope of constrained denoising.
The reason why the problems ($\mathcal{VCB}$), ($\mathcal{DCB}$), and ($\mathcal{GCB}$) are difficult to analyze is that distributional constraints are non-convex in terms of the denoiser. However, it is one of the fundamental insights of the theory of optimal transport (OT) (Villani, 2009) that problems involving such non-convex distributional constraints are in many settings equivalent to a suitable linear program over a space of couplings, and that analysis of the latter provides theoretical and computational advantages for the former. This connection has been previously made by García-Trillos and Sen (2024) at the population-level, and the present work extends this paradigm to the EB setting. This allows us to comprehensively study all the problems above, and to derive new results even for problem ($\mathcal{VCB}$) which has been extensively studied in previous literature.
1.3 Summary of Results
The main results of this paper concern the various constrained Bayes problems above, their EB approximations, extensions, and applications. We now describe these points in more detail.
Variance-Constrained Denoising. First, we provide a detailed study of problem ($\mathcal{VCB}$), which has been introduced in Louis (1984) and studied in many subsequent works (e.g., Ghosh (1992); Ghosh and Maiti (1999)). At the population level, we provide a closed-form solution for the variance-constrained Bayes denoiser $\delta_{\mathcal{VCB}}$ (Theorem 3.3) which shows that $\delta_{\mathcal{VCB}}$ is an affine function of the Bayes denoiser $\delta_{\mathcal{B}}$; crucially, this characterization arises by casting the denoising problem in terms of the geometry of the Bures-Wasserstein space of Gaussian OT, for which many explicit formulas are known. We further show that the variance-constrained Bayes risk $R_{\mathcal{VCB}}$ (the infimum of ($\mathcal{VCB}$)) is at most twice the optimal Bayes risk (the infimum of ($\mathcal{B}$)), illustrating that the price one needs to pay for matching the first two moments of the denoisers to that of the unknown latent variables is modest. At the empirical level, we provide an EB denoising procedure (Algorithm 1) which computes a denoiser $\hat{\delta}_{\mathcal{VCB}}$ whose distance to the oracle $\delta_{\mathcal{VCB}}$ exhibits an explicit rate of convergence (Theorem 4.3). See Figure 1.1 (third column) for an illustration of these variance-constrained denoisers in the simulation, which is consistent with them being affine functions of the corresponding unconstrained denoisers.
Distribution-Constrained Denoising. Next, we consider the distribution-constrained Bayes problem ($\mathcal{DCB}$). At the population level, we give conditions under which there is a uniquely defined distribution-constrained Bayes denoiser $\delta_{\mathcal{DCB}}$ (Theorem 3.8); at the empirical level, we provide an EB denoising scheme (Algorithm 2) which produces a denoiser $\hat{\delta}_{\mathcal{DCB}}$ that exhibits an explicit rate of convergence to the oracle constrained denoiser (Theorem 4.4). This work extends recent results in García-Trillos and Sen (2024) concerning the population-level problem; we provide rigorous statistical guarantees for an EB procedure which was stated (but not rigorously studied) in García-Trillos and Sen (2024, Appendix E). In Figure 1.1 (fourth column) we display the resulting denoisers in the simulation. The figure also highlights the difference between the problem of distribution-constrained EB denoising and the problem of deconvolution (Meister, 2009), that denoising implies a correspondence between the denoised data and the observations, while deconvolution does not.
General-Constrained Denoising. Third, we consider the general-constrained Bayes problem ($\mathcal{GCB}$), which encapsulates a wide variety of practical constraints (for instance, the fourth-moment constraints of Armstrong et al. (2022) or support constraints such as nonnegativity). While the precise statements and discussions of these results can be found in Appendix C, we briefly summarize them here. At the population level, we give conditions under which there is a uniquely-defined general-constrained Bayes denoiser $\delta_{\mathcal{GCB}}$; at the empirical level, we provide an EB denoising scheme which produces a denoiser $\hat{\delta}_{\mathcal{GCB}}$ that exhibits an explicit rate of convergence to the oracle constrained denoiser.
A fundamental flexibility of our paradigm is that it allows one to transform an arbitrary unconstrained EB denoiser into a suitable constrained EB denoiser, and the constrained denoiser inherits its rate of convergence from the unconstrained EB denoiser. For example, the unconstrained denoiser may arise via $G$-modeling, $F$-modeling, parametric EB, smoothed versions thereof, hybrids thereof, and more. This allows us to determine the rates of convergence for our constrained EB denoising procedures in many different settings of interest (e.g., light- and heavy-tailed $G$, Gaussian and Poisson likelihood $\{P_{\theta}\}_{\theta}$, conjugate parametric models), and we give a detailed analysis of the resulting rates of convergence in Appendix B. A notable example is the case of a nonparametric latent variable distribution and a Gaussian likelihood where we use nonparametric maximum likelihood (NPMLE) for $G$-modeling; our results imply that the rate of convergence of $\hat{\delta}_{\mathcal{VCB}}$ is nearly-parametric, and the rate of convergence of $\hat{\delta}_{\mathcal{DCB}}$ is slow (due to the slow minimax rate of nonparametric deconvolution, i.e., Carroll and Hall (1988); Zhang (1990); Fan (1991)).
We also illustrate our methodology in two applications (Section 6). In an application to astronomy (Section 6.1), we use our results to denoise the distribution of latent chemical abundances in a large catalog of observed stars, furthering the analyses of Soloff et al. (2024); Ratcliffe et al. (2020). In an application to baseball (Section 6.2), we use our results to denoise the distribution of latent batting skill for players who are promoted from minor- to major-league play. Because these applications involve some degree of heterogeneity in the observations (e.g., heteroskedasticity in the Gaussian case) we include a detailed discussion (Section D) of some considerations for extensions to the heterogeneous case.
1.4 Related Literature
We build upon a long tradition in the EB literature (Robbins, 1956; Efron, 2019) of constructing nonparametric denoisers that can provably perform nearly as well as oracle denoisers that operate under knowledge of the true latent variable distribution. The statistical properties of unconstrained denoisers that mimic the Bayes denoisers in ($\mathcal{B}$) have been increasingly well understood in recent years (Zhang, 1997; Brown and Greenshtein, 2009; Jiang and Zhang, 2009; Efron, 2011; Saha and Guntuboyina, 2020; Polyanskiy and Wu, 2021; Shen and Wu, 2022; Ignatiadis et al., 2023; Barbehenn and Zhao, 2023; Soloff et al., 2024; Ghosh et al., 2025). Here we lift existing results on unconstrained denoisers and provide further theory on EB approximations to constrained denoisers.
The fallacy of applying downstream statistical analyses to EB-denoised data sets has been recognized since the early developments of EB in astronomy. To wit, Sir Arthur Eddington introduced what is now known as Tweedie’s formula (Efron, 2011) to denoise noisy parallax observations; this method was communicated by Sir Frank Dyson (1926). Later, Eddington (1940) lamented that his formula “is not infrequently employed in an illegitimate way”: Astronomers would use the empirical distribution of denoised estimates $\widehat{\Theta}_{1},\dotsc,\widehat{\Theta}_{n}$ and interpret it as an improved estimate of the distribution of $\Theta_{1},\dotsc,\Theta_{n}$. Eddington explains that this practice is “fallacious,” since the distribution of $\widehat{\Theta}_{i}$ “deviates as far from the true distribution [of $\Theta_{i}$] in the direction of reduced spread as the observed distribution [of $Z_{i}$] does in the direction of increased spread.” Despite Eddington’s paper (which is well-known in astronomy), astronomers still often use shrunk values in the way Eddington warned against (Loredo, 2007).
The variance-constrained denoiser we study, traditionally called the constrained Bayes estimator, has been examined and used in applications by several authors (Louis, 1984; Lahiri, 1990; Ghosh, 1992; Devine and Louis, 1994; Ghosh and Maiti, 1999; Frey and Cressie, 2003; Bloom et al., 2017). While some results like Ghosh (1992, Theorem 1) apply generally to any Bayes estimator (paralleling our Theorem 3.3), existing applications predominantly pair constrained Bayes denoisers with parametric EB methods. To the best of our knowledge, only few authors (Shen and Louis, 2000; Lockwood et al., 2018; Lee et al., 2024) have explored more flexible nonparametric specifications alongside constrained Bayes denoising. Our work advances this literature in two ways: first, by deriving explicit estimation error rates that directly inherit from the original Bayes denoiser’s error rates; second, by building on OT theory to provide a new perspective and to extend variance constrained denoisers to multidimensional parameters. Though Ghosh and Maiti (1999) previously tackled multidimensional extensions, we identify and correct an oversight in their approach (see Remark 3.4).
Our proposed methodology is closely tied to practical considerations in a number of application areas. In this paper, we specifically consider applications to astronomy (Section 6.1) and baseball (Section 6.2). Regarding astronomy, Loredo (2007) has argued that distributional constraints on denoising procedures are a fundamental feature of astrostatistics. Another application is to baseball statistics; while our specific application is novel, it complements a long tradition of EB methods in baseball (Efron and Morris, 1975; Brown, 2008; Jiang and Zhang, 2010; Gu and Koenker, 2017a). Our methodology is directly relevant to domains that have recognized the need for constrained denoising and have employed variance-constrained denoisers, including multisite experiments in education research (Raudenbush and Bloom, 2015; Bloom et al., 2017; Lee et al., 2024), small-area estimation (Devine et al., 1994; Ghosh and Rao, 1994; Leyland and Davies, 2005), epidemiology (Lyles et al., 1997; Moore et al., 2010), and errors-in-variables regression (Freedman et al., 2004). We also anticipate applications in other fields, such as those in labor economics reviewed by Walters (2024).
2 Preliminaries on Optimal Transport
We review some aspects of optimal transport (OT), which provides the mathematical foundation for our main results. While OT was initially studied as a concrete problem in applied mathematics (Villani, 2009; Rachev and Rüschendorf, 1998), there has been an explosion of recent interest in applications to statistics (Panaretos and Zemel, 2020; Chewi et al., 2025).
We begin with Monge’s formulation of the OT problem, which aims to send each point to a given location in order to transport some source measure onto some target measure in such a way that minimizes the average cost. (For simplicity, we consider the case of quadratic cost function and of probability measures on the same space, but these can both be relaxed.) That is, for probability measures $\mu,\mu^{\prime}$ on $\mathbb{R}^{k}$, we consider the problem
$\underset{T:\mathbb{R}^{k}\to\mathbb{R}^{k}}{\textnormal{minimize}}\int_{\mathbb{R}^{k}}\|x-T(x)\|^{2}\textnormal{d}\mu(x)\qquad\textnormal{s.t.}\qquad T_{\#}\mu=\mu^{\prime}.$ (2.1)
Here $T_{\#}\mu$ denotes the pushforward of $\mu$ by $T$, defined as the probability measure on $\mathbb{R}^{k}$ given by the distribution of $T(X)$ when $X\sim\mu$. Problem (2.1) is difficult to solve directly, since the constraint $T_{\#}\mu=\mu^{\prime}$ is non-convex.
Another perspective is Kantorovich’s formulation, which aims to send each point to a distribution over locations in order to transport some source measure onto some target measure in a way that minimizes the average cost. That is, for probability measures $\mu,\mu^{\prime}$ on $\mathbb{R}^{k}$, we consider
$\underset{\pi\in\Gamma(\mu;\mu^{\prime})}{\textnormal{minimize}}\int_{\mathbb{R}^{k}}\|x-x^{\prime}\|^{2}\textnormal{d}\pi(x,x^{\prime})$ (2.2)
Here, $\Gamma(\mu;\mu^{\prime})$ denotes the set of all couplings of $\mu,\mu^{\prime}$, meaning the set of all joint distributions $\pi$ such that if $(X,X^{\prime})\sim\pi$ then $X\sim\mu$ and $X^{\prime}\sim\mu^{\prime}$. Problem (2.2) is a linear program (possibly infinite dimensional) so it is often simple to analyze and implement.
One of the fundamental insights of OT is that problems (2.1) and (2.2) are closely related. For instance, it is easy to show that (2.2) is a convex relaxation of (2.1) in the sense that it is a convex optimization problem whose domain naturally contains the domain of the former and its optimal value lower bounds the optimal value of the former. Remarkably, a result of Brenier (1991) provides sufficient conditions for (2.2) to be a tight convex relaxation of (2.1) in the sense that their optimal values agree, and every solution of the latter can be naturally mapped onto a solution of the former. More precisely, if $\mu$ and $\mu^{\prime}$ both have finite second moment and if $\mu$ has a density with respect to Lebesgue measure, then (2.1) admits a solution $\delta$ which is unique up to $\mu$-equivalence, (2.2) admits a unique solution $\pi$, and these solutions are related via $\pi=(\textnormal{id},\delta)_{\#}\mu$; moreover, the optimal $\delta$ for (2.1) must be equal to the gradient of some real-valued convex function defined on $\mathbb{R}^{m}$. In other words, the unique solution to the Kantorovich problem is a coupling supported on the graph of a gradient of a convex function which is a solution to the Monge problem.
In practice, a common approach to solving (2.1) is as follows. First, find a solution $\pi$ to (2.2). Second, define the function $\delta:\mathbb{R}^{k}\to\mathbb{R}^{k}$ via $\delta(x^{\prime}):=\int_{\mathbb{R}^{k}}x\,\pi(\textnormal{d}x,x^{\prime})\big{/}\int_{\mathbb{R}^{k}}\pi(\textnormal{d}x,x^{\prime})$ called the barycentric projection of $\pi$; note that this is just the conditional expectation $\mathbb{E}[X\,|\,X^{\prime}=x^{\prime}]$ when $(X,X^{\prime})\sim\pi$. If the convex relaxation above is tight, then $\delta$ must be solution to (2.1). Otherwise, $\delta$ can be thought of as an approximate solution to (2.1).
We next review some aspects of the geometry of the Wasserstein space. We write $\mathcal{P}_{2}(\mathbb{R}^{k})$ for the space of probability measures $\mu$ on $\mathbb{R}^{k}$ satisfying $\int_{\mathbb{R}^{k}}\|x\|^{2}\textnormal{d}\mu(x)<\infty$. For any probability measures $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{k})$, we write $W_{2}^{2}(\mu,\mu^{\prime})$ for the optimal value of problem (2.2) above, which is called the (square of the) 2-Wasserstein distance between $\mu$ and $\mu^{\prime}$. (If $\mu$ additionally has a density with respect to Lebesgue measure, then we know that this is equal to the optimal value of problem (2.1).) The metric space $(\mathcal{P}_{2}(\mathbb{R}^{k}),W_{2})$ is called the 2-Wasserstein space.
These notions simplify significantly in the case of centered Gaussian measures, which will prove useful in our study of the variance-constrained denoising problem ($\mathcal{VCB}$). For $m\in\mathbb{N}$, we write $\mathbb{K}(m)$ for the space of (strictly) positive definite $m\times m$ matrices. For any $\Sigma,\Sigma^{\prime}\in\mathbb{K}(m)$, we may consider problem (2.1) for probability measures $\mu=\mathcal{N}(0,\Sigma)$ and $\mu^{\prime}=\mathcal{N}(0,\Sigma^{\prime})$; it turns out that this problem has a (necessarily unique) solution given by the linear map $\delta(x):=\mathbf{t}_{\Sigma}^{\Sigma^{\prime}}x$, where $\mathbf{t}_{\Sigma}^{\Sigma^{\prime}}:=\Sigma^{-\nicefrac{{1}}{{2}}}(\Sigma^{\nicefrac{{1}}{{2}}}\Sigma^{\prime}\Sigma^{\nicefrac{{1}}{{2}}})^{\nicefrac{{1}}{{2}}}{\Sigma}^{-\nicefrac{{1}}{{2}}}$. Here, we write $A^{\nicefrac{{1}}{{2}}}$ for the unique positive definite square root of a positive definite matrix $A\in\mathbb{K}(m)$. We also write $\preceq$ and $\prec$ for the usual Loewner order on $\mathbb{K}(m)$, that is, the order generated by the positive semi-definite (PSD) cone. The space $\mathbb{K}(m)$ endowed with the metric inherited from $W_{2}$, under the natural identification, is referred to as the Bures-Wasserstein space.
3 Oracle Constrained Bayes Denoising
In this section, we consider the oracle versions (i.e., assuming that the latent variable distribution $G$ is known) of our constrained denoising problems ($\mathcal{VCB}$), ($\mathcal{DCB}$) and ($\mathcal{GCB}$). We address the case of variance constraints in Section 3.1, the case of distributional constraints in Section 3.2, and the case of general constraints in Section 3.3. The proofs of all the results stated in this section are given in Appendix F.
3.1 Variance-Constrained Denoising
We begin with a detailed study of problem ($\mathcal{VCB}$) in which our results can be made rather explicit. To give context for this problem, we mention the following motivating applications.
Example 3.1 (Mitigating underdispersion by “unshrinking” a bit).
Graphical distributional summaries (e.g., histograms, density plots) of Bayes denoisers inevitably exhibit underdispersion relative to the true distribution of the unobserved latent variables $\Theta_{1},\dotsc,\Theta_{n}$ (as directly follows from (1.2)). At an intuitive level, these graphical summaries would more faithfully represent the true heterogeneity in the population if the denoisers could be “unshrunk” to match the variance of the latent variables. Problem ($\mathcal{VCB}$) captures precisely this intuition through its variance-matching constraint, as Theorem 3.3 below demonstrates. Thus, problem ($\mathcal{VCB}$) provides a principled and useful strategy for presenting more faithful empirical summaries of heterogeneity by at least capturing the correct second moment structure. We note that this motivation was central to the work of Louis (1984) and Ghosh (1992).
Example 3.2 (Errors-in-variables regression: moment reconstruction).
Consider a regression setting with covariates $X\in\mathbb{R}^{k}$ and response $Y\in\mathbb{R}$; interest focuses on the relationship between $X$ and $Y$. We observe the covariates with noise, that is, we observe $W=X+E$ where $E$ has a known distribution. In regression calibration (Carroll and Stefanski, 1990) we perform the regression analysis on the pairs $(\mathbb{E}[X\mid W],Y)$; see Chen et al. (2025) for a critical discussion of the approach. By contrast, in moment reconstruction (Freedman et al., 2004), we seek a denoiser of $X$, which may possibly depend on both $W$ and $Y$, that satisfies
$\mathbb{E}[\delta(W,Y)]=\mathbb{E}[X],\qquad\textnormal{Cov}(\delta(W,Y))=\textnormal{Cov}(X),\qquad\mathbb{E}[\delta(W,Y)Y]=\mathbb{E}[XY],$
and apply the regression analysis on pairs $(\delta(W,Y),Y)$. The idea is that we would like our denoised covariates to preserve cross-moments with the response as these are fundamental for understanding the $X,Y$ relationship. A moment reconstruction denoiser is provided by optimization problem ($\mathcal{VCB}$) by taking $Z=(W,Y)$ and $\Theta=(X,Y)$.
Our analysis of problem ($\mathcal{VCB}$) requires a few assumptions, which we now introduce and discuss. Because we will impose covariance constraints on the denoising problem, we need some conditions to ensure that the constrained denoising problem admits a solution. The following is necessary:
$\int_{\mathbb{R}^{m}}\|\theta\|^{2}\textnormal{d}G(\theta)<\infty.$ (2M)
We require the second moment condition for $G$ but not for $F$ (the marginal distribution of $Z$). We also make the following assumption on the covariance of the Bayes denoiser $\delta_{\mathcal{B}}$:
$\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\textnormal{ is strictly positive definite}.$ (PD)
This condition is usually easy to verify; for instance, if $\Theta$ has Gaussian distribution $G=\mathcal{N}(\mu,A)$ for $A\succ 0$ and $P_{\theta}=\mathcal{N}(\theta,\Sigma)$, then we have $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))=A(A+\Sigma)^{-1}A\succ 0$ for any $\Sigma\succeq 0$. In this form, the assumption precludes the case that $\Theta$ is concentrated on a proper subspace of $\mathbb{R}^{m}$
Some additional assumptions provide sufficient regularity in order to ensure that the variance-constrained denoising problem admits a unique solution. For example, we assume
[…]
While condition (ZC) holds whenever all elements of the model $\{P_{\theta}\}_{\theta}$ have a density with respect to Lebesgue measure (for example, in the Gaussian case), it does not hold in some discrete cases (for example, in the Poisson case). Under assumption (ZC), assumption (BC) is quite mild (see García-Trillos and Sen (2024, Remark 2.2)).
We are ready to state the first of our main results, for which the univariate case is effectively due to Louis (1984); Ghosh (1992). Here, $R_{\mathcal{B}}$ denotes the minimum value of problem ($\mathcal{B}$) and $R_{\mathcal{VCB}}$ the minimum value of problem ($\mathcal{VCB}$). Also, recall the definition of $\mathbf{t}$ given in Section 2.
Theorem 3.3.
Under assumption (2M) and (PD), problem ($\mathcal{VCB}$) has solution
$\delta_{\mathcal{VCB}}(z)=\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}(\delta_{\mathcal{B}}(z)-\mathbb{E}[\Theta])+\mathbb{E}[\Theta].$ (3.1)
and its risk is given by $R_{\mathcal{VCB}}=R_{\mathcal{B}}+W_{2}^{2}(\mathcal{N}(0,\textnormal{Cov}(\Theta)),\mathcal{N}(0,\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))$. Under the further assumptions (ZC) and (BC), $\delta_{\mathcal{VCB}}$ is the $F$-a.s. unique solution to problem ($\mathcal{VCB}$).
The interpretation of the result is simple: The variance-constrained Bayes denoiser $\delta_{\mathcal{VCB}}$ is an affine function of the unconstrained Bayes denoiser $\delta_{\mathcal{B}}$, and the requisite rescaling is exactly the one that minimizes the OT cost from the covariance of the unconstrained Bayes denoiser $\delta_{\mathcal{B}}(Z)$ to the covariance of the latent variable $\Theta$.
Remark 3.4 (Comparison with Ghosh and Maiti (1999)).
The multivariate problem of optimal denoising subject to constraints on the covariance matrix has been previously studied by Ghosh and Maiti (1999). Their proposed denoiser is $\delta(z)=(\textnormal{Cov}(\Theta))^{\nicefrac{{1}}{{2}}}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))^{-\nicefrac{{1}}{{2}}}(\delta_{\mathcal{B}}(z)-\mathbb{E}[\Theta])+\mathbb{E}[\Theta]$. Note that this differs from the optimal denoiser given in Theorem 3.3, but both are of the form
$\delta(z)=T(\delta_{\mathcal{B}}(z)-\mathbb{E}[\Theta])+\mathbb{E}[\Theta]$ (3.2)
for a suitable matrix $T\in\mathbb{R}^{m\times m}$; that is, both are an affine transformation of the Bayes denoiser. We reconcile this difference by observing that an additional condition is needed in the discussion following Ghosh and Maiti (1999, equation (40)): Ghosh and Maiti (1999) identify that an optimal denoiser must be of the form (3.2), but omit the condition that $T$ be symmetric. Since all subsequent applications (e.g., to the dataset in Devine et al. (1994)) involve matrices $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$ that commute, it holds that $\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}=(\textnormal{Cov}(\Theta))^{\nicefrac{{1}}{{2}}}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))^{-\nicefrac{{1}}{{2}}},$ and so, in this case, the denoiser in Ghosh and Maiti (1999) is identical to the denoiser of Theorem 3.3. Stating the result in the general case when $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$ do not commute requires the machinery of the Bures-Wasserstein geometry as summarized in Section 2.
At this point, a natural goal is to quantify the statistical cost of imposing the moment matching constraint. More precisely, one may ask how large is the excess risk $R_{\mathcal{VCB}}-R_{\mathcal{B}}$. Our next result (proved in Appendix F) shows that the price to pay is modest.
Corollary 3.5.
We have $R_{\mathcal{B}}\leq R_{\mathcal{VCB}}\leq 2R_{\mathcal{B}}$.
3.2 Distribution-Constrained Denoising
Next we study the distribution-constrained Bayes problem ($\mathcal{DCB}$). This has recently been studied in García-Trillos and Sen (2024), and in this subsection we review some of their results and state them in a form that is amenable to our subsequent analyses. We keep in mind the following motivations throughout this subsection.
Example 3.6 (Distortion-perception tradeoff).
The problem of image restoration in computer vision involves a ground truth image $\Theta\in\mathbb{R}^{m}$ and an observed image $Z\in\mathbb{R}^{d}$, where the observation has much lower resolution compared to the ground truth or has been corrupted with noise. Blau and Michaeli (2018) refer to $\|\delta(Z)-\Theta\|^{2}$ as distortion and $\rho((\delta_{\mathcal{B}})_{\#}F,G)$ as perception for some metric or divergence $\rho$ on the space of probability measures, and they show that simultaneous minimization of distortion and perception is not possible. Freirich et al. (2021) proposed to take $\rho=W_{2}$ and to try to minimize the distortion subject to a constraint on the perception, which corresponds to problem ($\mathcal{DCB}$) when we impose the constraint of perfect perception.
Example 3.7 (Errors-in-variables regression continued).
In Example 3.2, we introduced moment reconstruction. Freedman et al. (2004) propose moment reconstruction with the following motivation. Suppose we can come up with a denoiser $\delta(W,Y)$ of $X$ such that, $(\delta(W,Y),Y)\overset{\mathcal{D}}{=}(X,Y),$ then we could learn about the relationship between $(X,Y)$ by studying $(\delta(W,Y),Y)$. Note that the constraint above corresponds to the distribution-constrained Bayes problem in ($\mathcal{DCB}$). Freedman et al. (2004) instead propose the moment reconstruction approach described in Example 3.2, arguing that enforcing the full distributional constraint “is a very difficult problem, but if we content ourselves with the lesser aim of matching just the first two moments of the joint distribution, then a simple solution can be obtained.”
Our main result in this subsection is Theorem 3.8 which says that problem ($\mathcal{DCB}$) admits a tight convex relaxation. To state this, we recall that $F(\textnormal{d}z)=\int_{\mathbb{R}^{m}}P_{\theta}(\textnormal{d}z)\textnormal{d}G(\theta)$ is the marginal distribution of $Z$, and let us define the (non-standard) cost function $c_{G}:\mathbb{R}^{d}\times\mathbb{R}^{m}\to[0,\infty)$
$c_{G}(z,\eta):=\|\eta-\delta_{\mathcal{B}}(z)\|^{2}$ (3.3)
which, as we emphasize in the notation, depends on the unknown distribution $G$ (and also the known likelihood $P=\{P_{\theta}\}_{\theta}$) via the conditional expectation $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$. We then have the following, which is a re-statement of García-Trillos and Sen (2024, Theorem 2.4) and bears some similarity to Freirich et al. (2021, Theorem 1). Recall from Section 2 that $\Gamma(F;G)$ represents the space of all couplings of $F$ and $G$, and let us write $R_{\mathcal{DCB}}$ for the minimum value of problem ($\mathcal{DCB}$).
Theorem 3.8.
Under assumptions (2M), (ZC), and (BC), the problem
$\underset{\pi\in\Gamma(F;G)}{\textnormal{minimize}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\,\textnormal{d}\pi(z,\eta),$ (3.4)
has a unique solution $\pi_{\mathcal{DCB}}$. This solution is concentrated on the graph of a function which we denote by $\delta_{\mathcal{DCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. This function $\delta_{\mathcal{DCB}}$ can be written as $\delta_{\mathcal{DCB}}=\nabla\phi\circ\delta_{\mathcal{B}}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, and we have $R_{\mathcal{DCB}}=R_{\mathcal{B}}+W_{2}^{2}(G,(\delta_{\mathcal{B}})_{\#}F)$. Consequently, (3.4) is a tight convex relaxation of ($\mathcal{DCB}$), and $\delta_{\mathcal{DCB}}$ is the $F$-a.s. unique solution to ($\mathcal{DCB}$).
The above result says that, under appropriate conditions, problem ($\mathcal{DCB}$) indeed has a unique solution which can be obtained by solving the linear program (3.4). Further, the risk of the distribution-constrained denoiser $\delta_{\mathcal{DCB}}(Z)$ is the sum of the Bayes risk $R_{\mathcal{B}}$ and the squared Wasserstein distance between the distributions of $\Theta$ and $\delta_{\mathcal{B}}(Z)$. Contrast this with Theorem 3.3, where we showed that the risk of the variance-constrained denoiser $\delta_{\mathcal{VCB}}(Z)$ is sum of the Bayes risk and the squared Bures-Wasserstein distance between the covariance matrices of $\Theta$ and $\delta_{\mathcal{B}}(Z)$.
3.3 General-Constrained Denoising
Lastly, we consider problem ($\mathcal{GCB}$) of general-constrained Bayes denoising, which allows enough flexibility to encapsulate problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$) as well as many others. Appendix C.1 contains our detailed discussion, including conditions under which problem ($\mathcal{GCB}$) admits a tight convex relaxation and a unique solution $\delta_{\mathcal{GCB}}$ (Theorem C.3), which is similar to the analogous result for problem ($\mathcal{DCB}$) (Theorem 3.8). As an example of a general-constrained denoising problem relevant to our application in Subsection 6.2 (see Appendix C for more examples), we consider the following:
Example 3.9 (Variance constraint and support constraint).
Suppose that $G$ is known to have support contained in a closed set $S\subseteq\mathbb{R}^{m}$. In that case, the posterior mean satisfies $\delta_{\mathcal{B}}(Z)\in S$ almost surely. However, the variance constrained denoiser $\delta_{\mathcal{VCB}}(\cdot)$ may take on values in $S$ with positive probability, which can be undesirable. Instead, one may aim to minimize the denoising risk subject to simultaneous constraints on the variance and the support of $\delta(Z)$, thereby solving
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\,\,\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\quad\textnormal{s.t.}\quad\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta],\,\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta),\,\,\mathbb{P}(\delta(Z)\in S)=1.$ (3.5)
To see that this can be cast as a special case of problem ($\mathcal{GCB}$), take $\psi(\eta):=\min_{\theta\in S}\|\eta-\theta\|$ and note that $\mathbb{P}(\delta(Z)\in S)=1$ is equivalent to $\mathbb{E}[\psi(\delta(Z))]=\mathbb{E}[\psi(\Theta)]=0$ In the case $m=d=1$ an important constraint on the support is nonnegativity, i.e., $S=[0,\infty)$.
4 Empirical Constrained Bayes Denoising
In this section, we consider constrained denoising at the empirical level, i.e., we only observe $Z_{1},\ldots,Z_{n}$ from the independent, identically-distributed (i.i.d.) sequence of pairs $(\Theta_{1},Z_{1}),\ldots,(\Theta_{n},Z_{n})$ coming from model (1.1). Importantly, we assume throughout this section that the distribution $G$ of $\Theta_{1},\ldots,\Theta_{n}$ is unknown; it must therefore be estimated (either implicitly or explicitly) in order to apply the results of Section 3 about the corresponding oracle constrained Bayes denoising problems.
Specifically, we suppose that $\delta_{\ast}$ is one of $\delta_{\mathcal{VCB}},\delta_{\mathcal{DCB}}$, or $\delta_{\mathcal{GCB}}$, and we aim to develop an EB estimator $\hat{\delta}_{\ast}$ of $\delta_{\ast}$. All of our estimators arise by applying some further processing to a denoiser $\hat{\delta}_{\mathcal{B}}$ which is already known to be a good approximation of the Bayes optimal (that is, unconstrained) denoiser $\delta_{\mathcal{B}}$. More precisely, if we assume a priori that $\hat{\delta}_{\mathcal{B}}$ satisfies $\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|^{2}\overset{\mathbb{P}}{\to}0$, then we can construct $\hat{\delta}_{\ast}$ from $\hat{\delta}_{\mathcal{B}}$ in such a way that we have $\hat{\delta}_{\ast}\to\delta_{\ast}$ as $n\to\infty$; moreover the rate of convergence of $\hat{\delta}_{\ast}$ to $\delta_{\ast}$ depends explicitly on the rate of convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$. This flexibility in choosing $\hat{\delta}_{\mathcal{B}}$ is important for the broad applicability of the methodology; the reader can keep in mind the following examples for how EB denoisers may arise.
Example 4.1 ($G$-modeling).
If it is possible to derive an estimator $\hat{G}$ that suitably converges to the unknown $G$, then an EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$ arises by computing the posterior mean of $\Theta$ given $Z$, assuming that $\Theta\sim\hat{G}$. Typically, $\hat{G}$ arises by solving the following nonparametric maximum likelihood estimation (NPMLE) problem
$\hat{G}_{n}\;\in\;\underset{H\in\mathcal{P}(\mathbb{R}^{m})}{\arg\max}\,\frac{1}{n}\sum_{i=1}^{n}\log\left(\int_{\mathbb{R}^{m}}p_{\theta}(Z_{i})\textnormal{d}H(\theta)\right),$ (4.1)
where $p_{\theta}$ is the density of $P_{\theta}$ with respect to some fixed reference measure (usually the Lebesgue measure or the counting measure) for each $\theta$. The main applications of $G$-modeling are the setting where $\{P_{\theta}\}_{\theta}$ is a Gaussian location family studied by Jiang and Zhang (2009); Saha and Guntuboyina (2020); Soloff et al. (2024) and the setting where $\{P_{\theta}\}_{\theta}$ is the Poisson family studied by Shen and Wu (2022); Polyanskiy and Wu (2021); these cited works show that the convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$ achieves nearly the parametric rate, that is $n^{-1}$ up to logarithmic factors, whenever $G$ is light-tailed. This partially explains the remarkable success of EB: while deconvolution is hard, estimation of the Bayes denoiser is easy.
Example 4.2 (Smooth $G$-modeling).
If it is known that the prior $G$ is itself a mixture of Gaussians whose components have variance satisfying some lower bound (in the positive semi-definite order), then one can restrict the optimization in (4.1) accordingly, leading to the so-called smooth NPMLE introduced by Magder and Zeger (1996). When such restrictions are well-specified (i.e., when the user-specified component variance is at least as small as the component variance in $G$), the rate of convergence is identical to that of the unrestricted NPMLE in Example 4.1. In some settings, the increased regularity of smooth NPMLE over generic NPMLE is desirable.
In addition to $G$-modeling and smooth $G$-modeling, we note that other EB approaches are possible, including $F$-modeling and conjugate parametric models. In Appendix A we describe these in more detail, including some explicit formulas in the case of conjugate parametric models. Our results in this section are divided into variance constraints (Subsection 4.1), distributional constraints (Subsection 4.2), and general constraints (Subsection 4.3). The proofs of all results in this section can be found in Appendix G.
4.1 Variance-Constrained Denoising
We begin with problem ($\mathcal{VCB}$) in which our results are the most complete. That is, we consider finding, on the basis of the observations $Z_{1},\ldots,Z_{n}$ and the known likelihood $\{P_{\theta}\}_{\theta}$, an estimator $\hat{\delta}_{\mathcal{VCB}}$ of the solution $\delta_{\mathcal{VCB}}$ to problem ($\mathcal{VCB}$), defined in (3.1). Note that $\delta_{\mathcal{VCB}}$ need not be the unique solution to ($\mathcal{VCB}$), but Theorem 3.3 guarantees that it always exists and hence it is an oracle denoiser that we may target.
From the form of $\delta_{\mathcal{VCB}}$ in (3.1), it is clear that, in addition to the existence of an EB denoiser $\hat{\delta}_{\mathcal{B}}$, we also need to estimate the matrices $\textnormal{Cov}(\Theta)$ and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))$. The latter can be estimated via the empirical covariance matrix of $\hat{\delta}_{\mathcal{B}}(Z_{1}),\ldots,\hat{\delta}_{\mathcal{B}}(Z_{n})$, and the former can be estimated from the data $Z_{1},\ldots,Z_{n}$ and the likelihood $\{P_{\theta}\}_{\theta}$ by applying the law of total covariance in a suitable way. More precisely, we propose a constrained denoising procedure in Algorithm 1 and our main result below will provide an explicit rate of convergence for the resulting constrained denoiser.
Let us describe the assumptions that will be used in our main result. First, we need some form of integrability of the prior $G$, which is similar to but slightly stronger than what we assumed in (2M) for the oracle problem. Namely, we assume
$\int_{\mathbb{R}^{m}}\|\theta\|^{q}\textnormal{d}G(\theta)<\infty\qquad\textnormal{for some}\qquad 2<q\leq 4,$ (FM)
and the exponent $q$ will manifest in the rate of convergence below.
Next, we need some assumptions that allow us to estimate $\textnormal{Cov}(\Theta)$. For example, we require the following, which states that the likelihood $\{P_{\theta}\}_{\theta}$ is parameterized by its mean:
$\mathbb{E}[Z\,|\,\Theta]=\Theta\qquad\textnormal{ almost surely.}$. (M)
This assumption forces $d=m$ and is satisfied for most of the usual likelihoods $\{P_{\theta}\}_{\theta}$, including the Gaussian and Poisson cases (but note that it forces us to use “‘scale parameterizations” rather than “rate parameterizations” in, e.g., geometric and exponential models). Additionally, we require the existence of an estimator $\hat{\Sigma}_{n}$ of $\Sigma:=\mathbb{E}[\textnormal{Cov}(Z\,|\,\Theta)]$ with a parametric rate of convergence
$\|\hat{\Sigma}_{n}-\Sigma\|^{2}=O_{\mathbb{P}}\left(n^{-1}\right).$ (EC)
This condition is satisfied in the Gaussian case since $\Sigma$ is exactly known and can be estimated via $\hat{\Sigma}_{n}:=\Sigma$, and it is satisfied in the Poisson case since (M) implies $\Sigma=\mathbb{E}[Z]$ and the latter can be estimated at rate $n^{-1}$ via the sample mean of $Z_{1},\ldots,Z_{n}$, by virtue of (FM) (or even (2M)). Since the law of total covariance yields
$\textnormal{Cov}(Z)=\mathbb{E}[\textnormal{Cov}(Z\,|\,\Theta)]+\textnormal{Cov}(\mathbb{E}[Z\,|\,\Theta])=\Sigma+\textnormal{Cov}(\Theta),$ (4.2)
we may use assumptions (M) and (EC) to estimate $\textnormal{Cov}(\Theta)$ via $(\hat{S}_{n}-\hat{\Sigma}_{n})_{+}$, where $(\,\cdot\,)_{+}$ denotes positive semi-definite truncation and $\hat{S}_{n}$ denotes the empirical covariance matrix of $Z_{1},\ldots,Z_{n}$.
Last, we assume that we are equipped with some EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$, and that it satisfies the following quantitative notion of convergence: For some specified sequence $\alpha_{n}\to 0$, we have
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}(\alpha_{n}),$ (EBQ)
and the rate $\alpha_{n}$ will manifest in the rate of convergence below. From the results cited in the examples at the beginning of this section, we may determine $\alpha_{n}$ in many models of interest.
Algorithm 1 An EB approximation $\hat{\delta}_{\mathcal{VCB}}$ of the oracle variance-constrained Bayes denoiser $\delta_{\mathcal{VCB}}$.
1:procedure VarianceConstrainedEB($Z_{1},\ldots,Z_{n};\{P_{\theta}\}_{\theta}$)
2:     input: samples $Z_{1},\ldots,Z_{n}\in\mathbb{R}^{m}$, likelihood $\{P_{\theta}\}_{\theta}$
3:     output: denoising function $\hat{\delta}_{\mathcal{VCB}}:\mathbb{R}^{m}\to\mathbb{R}^{m}$
4:     $\hat{\delta}_{\mathcal{B}}(\,\cdot\,)\leftarrow$ EB approximation of $\delta_{\mathcal{B}}(\,\cdot\,)$
5:     $\hat{\Sigma}_{n}\leftarrow$ estimate of $\Sigma$
6:     $\hat{M}\leftarrow$ sample covariance matrix of $\hat{\delta}_{\mathcal{B}}(Z_{1}),\ldots\hat{\delta}_{\mathcal{B}}(Z_{n})$
7:     $\hat{\mu}\leftarrow$ sample mean of $Z_{1},\ldots,Z_{n}$
8:     $\hat{S}\leftarrow$ sample covariance matrix of $Z_{1},\ldots,Z_{n}$
9:     $\hat{A}\leftarrow(\hat{S}-\hat{\Sigma}_{n})_{+}$
10:     $\hat{\mathbf{t}}\leftarrow\hat{M}^{-\nicefrac{{1}}{{2}}}(\hat{M}^{\nicefrac{{1}}{{2}}}\hat{A}\hat{M}^{\nicefrac{{1}}{{2}}})^{\nicefrac{{1}}{{2}}}\hat{M}^{-\nicefrac{{1}}{{2}}}$
11:     $\hat{\delta}_{\mathcal{VCB}}(\,\cdot\,)\leftarrow\hat{\mathbf{t}}(\hat{\delta}_{\mathcal{B}}(\,\cdot\,)-\hat{\mu})+\hat{\mu}$
12:     return $\hat{\delta}_{\mathcal{VCB}}$
13:end procedure
This leads us to the main result, whose proof appears in Appendix G.1:
Theorem 4.3.
Under (M), (FM), (EC), and (EBQ), the denoiser $\hat{\delta}_{\mathcal{VCB}}$ from Algorithm 1 satisfies
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})\right\|_{2}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)\qquad\textnormal{as}\qquad n\to\infty.$
The conclusion of Theorem 4.3 for the denoiser given in Algorithm 1 is that the rate of convergence of variance-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising (line 4) and covariance estimation (line 8) is slower. Moreover, Appendix B contains a detailed discussion of the rates of convergence in the Gaussian and Poisson models arising from this result, and we observe that it is possible for either of these terms to dominate. In the case of a Gaussian likelihood and a subexponential latent variable distribution, we note that the resulting rate is nearly parametric, i.e. $n^{-1}$ up to logarithmic factors. In the case of an an exponential family likelihood and a conjugate prior, the resulting rate is exactly parametric.
4.2 Distribution-Constrained Denoising
Next we consider the case of distributional constraints, as in problem ($\mathcal{DCB}$). Recall that in Subsection 3.2 we provided sufficient conditions for the existence of a unique solution $\delta_{\mathcal{DCB}}$ to problem ($\mathcal{DCB}$); our present goal is to develop an EB methodology which targets this denoiser.
From the characterization of $\delta_{\mathcal{DCB}}$ as the solution to the OT problem (3.4), it is clear that, in addition to the existence of an EB denoiser $\hat{\delta}_{\mathcal{B}}$, we also need to be able to estimate the distribution $F$, the distribution $G$, and the cost function $c_{G}$. (Recall its definition in (3.3).) We may estimate $F$ via the empirical distribution of $Z_{1},\ldots,Z_{n}$, $G$ as $\hat{G}_{n}$ via $G$-modeling (as in, e.g., Examples 4.1 and 4.2), and $c_{G}$ as a function of the given EB denoiser $\hat{\delta}_{\mathcal{B}}$; then we use these approximations in an OT problem paralleling (3.4), and the barycentric projection of its optimal coupling is an approximation of $\delta_{\mathcal{DCB}}$. (Recall from Section 2 that the barycentric projection recovers the OT map if the optimal coupling is supported on the graph of a function, and that it provides an approximation of the OT map otherwise.) We define this procedure explicitly in Algorithm 2, and we prove a rate of convergence as our main result below. Details for implementing Algorithm 2 in practice can be found in Appendix E.
Let us describe the assumptions needed in order to state and prove our main result. First, as described above, we need a consistent estimator $\hat{G}_{n}$ of $G$, meaning that for some specified sequence $\beta_{n}\to 0$ we have
$W_{2}^{2}(\hat{G}_{n},G)=O_{\mathbb{P}}(\beta_{n}).$ (D)
In most cases, we take $\hat{G}_{n}$ to be the NPMLE of $G$ (or a smoothed version thereof), in which case many recent results (e.g., Soloff et al. (2024); Polyanskiy and Wu (2021)) can be used to determine $\beta_{n}$. More generally, estimation of $G$ is referred to as deconvolution and it is known that optimal rates of convergence can be slow (Carroll and Hall, 1988; Fan, 1991; Zhang, 1990). Second, we need to quantify convergence of the empirical distribution of $\Theta_{1},\ldots,\Theta_{n}$ to $G$ in the Wasserstein metric; that is, for some sequence $\gamma_{n}\to 0$, we assume
$W_{2}^{2}\left(\bar{G}_{n},\;G\right)=O_{\mathbb{P}}(\gamma_{n})\qquad\textnormal{ where }\qquad\bar{G}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\Theta_{i}}.$ (EW)
It is known that $\gamma_{n}$ depends on the dimension $m$, and explicit forms of $\gamma_{n}$ can be found in Fournier and Guillin (2015) and Chewi et al. (2025, Section 2.9). Third, recall that, under assumptions (2M), (ZC), and (BC), Theorem 3.8 guarantees that problem ($\mathcal{DCB}$) admits a unique solution $\delta_{\mathcal{DCB}}=\nabla\phi\circ\delta_{\mathcal{B}}$; presently, we further assume the regularity that
$\phi\textnormal{ is $\lambda$-strongly convex and $L$-smooth for some }\lambda,L>0.$ (R)
While this condition may be hard to verify in practice, it is indeed commonly assumed in literature which proves rates of convergence for estimation of OT maps (e.g., Manole and Niles-Weed (2024); Slawski and Sen (2024); Deb et al. (2021)).
Algorithm 2 An EB estimator $\hat{\delta}_{\mathcal{DCB}}$ of the oracle distribution-constrained Bayes denoiser $\delta_{\mathcal{DCB}}$.
1:procedure DistributionConstrainedEB($Z_{1},\ldots,Z_{n};\{P_{\theta}\}_{\theta}$)
2:     input: samples $Z_{1},\ldots,Z_{n}$, likelihood $\{P_{\theta}\}_{\theta}$
3:     output: denoising function $\hat{\delta}_{\mathcal{DCB}}:\{Z_{1},\ldots,Z_{n}\}\to\mathbb{R}^{m}$
4:     $\hat{\delta}_{\mathcal{B}}(\,\cdot\,)\leftarrow$ EB approximation of $\delta_{\mathcal{B}}(\,\cdot\,)$
5:     $\hat{G}_{n}\leftarrow$ estimate of $G$
6:     $\hat{c}_{n}(Z_{i},\eta)\leftarrow\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\eta\|^{2}$ for all $1\leq i\leq n$ and $\eta\in\mathbb{R}^{m}$
7:     $\hat{\pi}_{\mathcal{DCB}}\leftarrow$ minimize $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}(z,\eta)\textnormal{d}\pi(z,\eta)$
8:                 over         probability measures $\pi\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{m})$
9:                 with        $\pi(\{Z_{i}\}\times\mathbb{R}^{m})=\frac{1}{n}$ for all $1\leq i\leq n$
10:                 and         $\pi(\mathbb{R}^{d}\times\textnormal{d}\eta)=\hat{G}_{n}(\textnormal{d}\eta)$
11:     $\hat{\delta}_{\mathcal{DCB}}(Z_{i})\leftarrow\int_{\mathbb{R}^{m}}\eta\,\textnormal{d}\hat{\pi}_{\mathcal{DCB}}(\eta\,|\,Z_{i})$ for all $1\leq i\leq n$
12:     return $\hat{\delta}_{\mathcal{DCB}}$
13:end procedure
These considerations lead us to our next main result.
Theorem 4.4.
Under assumptions (ZC), (BC), (R), (D), (EW), and (EBQ), the denoiser $\hat{\delta}_{\mathcal{DCB}}$ from Algorithm 2 satisfies
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{DCB}}(Z_{i})-\delta_{\mathcal{DCB}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}\left(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\beta_{n}\vee\gamma_{n}\right)\qquad\textnormal{as}\qquad n\to\infty.$ (4.3)
The conclusion of Theorem 4.4 for the denoiser given in Algorithm 2 is that the rate of convergence of distribution-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising (line 4), deconvolution (line 5), and convergence of the empirical Wasserstein distance is slowest. While in Appendix B.2 we give a detailed discussion of the resulting rates of convergence in some concrete settings of interest, we emphasize that $\beta_{n}$ typically dominates in nonparametric problems and that $\gamma_{n}$ typically dominates in parametric problems. The term $\alpha_{n}^{\nicefrac{{1}}{{2}}}$ is likely an artifact of the proof and we expect that it can be improved to $\alpha_{n}$ with further work, but we do not pursue this since it does not affect the rates in cases of interest.
4.3 General-Constrained Denoising
In Appendix C.2, we develop an EB methodology for targeting the oracle general-constrained denoiser $\delta_{\mathcal{GCB}}$. Methodologically, we regard ($\mathcal{GCB}$) as a way to tradeoff between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$); the former only encodes limited information about the latent variable distribution (the first two moments), the latter suffers from slow rates of convergence (due to the difficulty of the deconvolution problem) and may have larger denoising risk. Often some intermediate behavior may be desired. The rate of convergence of Theorem C.5 reflects the fact that ($\mathcal{GCB}$) can interpolate between ($\mathcal{VCB}$) and ($\mathcal{DCB}$).
5 Heterogeneous (Empirical) Constrained Bayes Denoising
Our results and methods so far have been developed under the assumption that all $n$ observations follow model (1.1). While this setting covers many important applications, it excludes settings with heterogeneity in the likelihood, e.g., settings with varying precision in the observations, as well as settings with heterogeneity in the prior. To accommodate such heterogeneity, we extend model (1.1) by including a heterogeneity parameter $\Xi\in\mathbb{R}^{p}$ in addition to the latent variable $\Theta$ and the observation $Z$:
$(\Theta\,|\,\Xi=\xi)\sim G_{\xi}\qquad\mbox{and}\qquad(Z\mid\Theta=\theta,\Xi=\xi)\sim P_{\theta,\xi}.$ (5.1)
Here, $\{G_{\xi}\}_{\xi}$ is an unknown family of probability measures on $\mathbb{R}^{m}$ and $\{P_{\theta,\xi}\}_{\theta,\xi}$ is a known family of probability distributions on $\mathbb{R}^{d}$. We assume that $\Theta$ is unobserved while $(Z,\Xi)$ is observed. We also consider the EB setting where we have i.i.d. samples $(\Theta_{1},\Xi_{1},Z_{1}),\ldots,(\Theta_{n},\Xi_{n},Z_{n})$ from (5.1) and we observe $(\Xi_{1},Z_{1}),\ldots,(\Xi_{n},Z_{n})$.
In our application to astronomy in Subsection 6.1, we model measurements as $Z_{i}\sim\mathcal{N}(\Theta_{i},\Sigma_{i})$ where the covariance matrices $\Sigma_{i}$ vary across observations. This fits framework (5.1) by setting $\Xi_{i}=\Sigma_{i}$ and $P_{\theta,\xi}=\mathcal{N}(\theta,\xi)$, allowing us to handle the heterogeneous measurement uncertainty in astronomical data. In our baseball application in Subsection 6.2, part of our modeling involves Poisson observations with heterogeneous exposure, where we set $\Xi_{i}=\lambda_{i}$ and $P_{\theta,\xi}=\textnormal{Poi}(\xi\theta)$. We call $\lambda_{i}$ the exposure (Norberg, 1979), representing the number of games played by player $i$, so that $\Theta_{i}=\mathbb{E}[Z_{i}/\lambda_{i}]$ is the rate normalized by exposure time. In both applications, we assume that $\{G_{\xi}\}_{\xi}$ does not depend on $\xi$, i.e., there is some unknown $G$ satisfying $G_{\xi}=G$ for all $\xi$ [^1] Allowing for dependence of $G_{\xi}$ on $\xi$ is important in some applications, see Appendix D.1 and Chen (2022).
In Appendix D, we discuss applications of (5.1) in more detail (Appendix D.1) and describe modifications of the EB procedures in Section 4 to construct constrained denoisers under model (5.1). Constraints on the denoising procedure have two possible forms, namely a constraint on the marginal distribution of $\delta(Z,\Xi)$ or a constraint on the conditional distribution of $\delta(Z,\Xi)$ given $\{\Xi=\xi\}$ for almost every $\xi$. In particular, we discuss marginal and conditional variance constraints (Appendix D.2) leading to denoisers $\delta_{\mathcal{M}\mathcal{VCB}}$ and $\delta_{\mathcal{C}\mathcal{VCB}}$ respectively, marginal distribution constraints (Appendix D.3) leading to a denoiser $\delta_{\mathcal{M}\mathcal{DCB}}$, and marginal general constraints (Appendix D.4), leading to a denoiser $\delta_{\mathcal{M}\mathcal{GCB}}$; as usual, we write $\hat{\delta}_{\ast}$ to denote an EB approximation of an oracle denoiser $\delta_{\ast}$. See also Section D.2.3 for a numerical illustration of the difference between the marginal conditional variance constraints.
6 Applications
In this section, we consider applications of our methodology to two important areas which have motivated much of the EB theory throughout its development.
6.1 Astronomy
First we consider an application to the problem of denoising the relative chemical abundances present in a catalog of stars. In short, the relative abundances of various chemicals influence stellar formation, evolution, and dynamics, hence their estimation is a fundamental task in astronomy.
Our data comes from the Apache Point Observatory Galactic Evolution Experiment survey (APOGEE); see Majewski et al. (2017); Abolfathi et al. (2018); Ratcliffe et al. (2020); Soloff et al. (2024) for further detail. More specifically, we focus on Oxygen-Iron and Nitrogen-Iron relative abundances (denoted [O/Fe] and [N/Fe], respectively) for stars in the red clump catalog, following the analysis of Ratcliffe et al. (2020). For each star ($n=2{,}000$ stars sampled at random from the full catalog of $2.7\times 10^{4}$ stars), we have an estimate of the values [O/Fe] and [N/Fe], and the measurement error of these estimates is known.
To state our model explicitly, we make the following modeling assumptions: 1. Every star $i$ has latent chemical abundances $\Theta_{i}=(\Theta_{i}^{\textnormal{[O/Fe]}},\Theta_{i}^{\textnormal{[N/Fe]}})$, and $\Theta_{1},\ldots,\Theta_{n}$ are i.i.d. from an unknown distribution $G$ on $\mathbb{R}^{2}$. 2. The measurement of every star $i$ admits error given by a diagonal covariance matrix $\Sigma_{i}$, the matrices $\Sigma_{1},\ldots,\Sigma_{n}$ are i.i.d. from some unknown distribution on $\mathbb{K}(2)$, and $\Theta_{1},\ldots,\Theta_{n}$ are independent of $\Sigma_{1},\ldots,\Sigma_{n}$. 3. For each star $i$, the observed pair of measurements $Z_{i}=(Z_{i}^{\textnormal{[O/Fe]}},Z_{i}^{\textnormal{[N/Fe]}})$ has conditional distribution $\mathcal{N}(\Theta_{i},\Sigma_{i})$ given $\Theta_{1},\ldots,\Theta_{n},\Sigma_{1},\ldots,\Sigma_{n}$, and $Z_{1},\ldots,Z_{n}$ are conditionally independent given $\Theta_{1},\ldots,\Theta_{n},\Sigma_{1},\ldots,\Sigma_{n}$. The assumptions above are common in existing literature, e.g., each $\Sigma_{i}$ being diagonal (Ting and Weinberg, 2022, page 16), and $\Theta_{i}$ and $\Sigma_{i}$ being independent (Soloff et al., 2024, Section 1.1).
Next we describe the denoising methods of interest. All of our approaches will be based on smooth $G$-modeling, as in Example 4.2, where we assume that $G$ is a Gaussian location mixture model, where the covariance matrix of each Gaussian component is bounded below, in the PSD order, by $\tau^{2}I_{2}$, where $\tau^{2}=0.0025$. Then, we compute the heteroskedastic multivariate NPMLE of Soloff et al. (2024) over this restricted class of $G$, yielding an unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$. Last, we use $\hat{\delta}_{\mathcal{B}}$ to compute the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal distribution-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{DCB}}$, as described in Section D.
Figure 6.1: Denoising the stellar relative chemical abundances dataset of Subsection 6.1. We plot the raw data (first), the unconstrained EB denoised data (second), the marginal variance-constrained EB denoised data (third), and the marginal distribution-constrained EB denoised data (fourth).
The results are shown in Figure 6.1, which deserves some discussion. First, note that the EB denoiser $\hat{\delta}_{\mathcal{B}}$ recovers some interesting latent low-dimensional structure from the data. However, the difference between the EB denoiser $\hat{\delta}_{\mathcal{B}}$ and the variance-constrained EB denoiser $\hat{\delta}_{\mathcal{VCB}}$ illustrates that the shrinkage of $\hat{\delta}_{\mathcal{B}}$ may yield overestimation of this latent structure. Also, we note that the transformation from the EB denoiser $\hat{\delta}_{\mathcal{B}}$ to the variance-constrained EB denoiser $\hat{\delta}_{\mathcal{VCB}}$ is an affine but it is not merely scaling; the upper branch slightly changes its angle, which suggests that the correlation between the two components may be underestimated by the EB denoiser. Lastly, we note that the distribution-constrained EB denoiser $\hat{\delta}_{\mathcal{DCB}}$ reveals a similar latent low-dimensional structure, but that the resulting distribution is more equispaced and this allows one identify more subtle features in the distribution (e.g., the separated cluster appearing at the ends of the top branch, and the positions of points outside the bulk of the data set).
6.2 Baseball
Second, we consider an application to denoising the joint distribution of minor-league batting skill and major-league batting skill for rookie baseball players. [^2] Strictly speaking, a “rookie” player is one who is in his first season meeting at least one criterion involving the number of games played or the number of at-bats. In this paper, we simply mean a player that is in his first major league appearance, since number of games played will be a part of our model. For the sake of simplicity, we focus on the statistic RBI (runs batted in) which is a common measure of batter performance.
Our data consists of all batters in their first season of major league play during the 2022 to 2024 season of the MLB which is publicly available through the website FanGraphs. [^3] Available at: https://www.fangraphs.com/ For each player ($n=324$), we observe his RBIs and his number of games played, in both their final minor league season and their initial major league season. To state our model explicitly, we make the following modeling assumptions: 1. Every player $i$ has a latent bivariate skill $\Theta_{i}=(\Theta_{i}^{\textnormal{min}},\Theta_{i}^{\textnormal{maj}})$, and $\Theta_{1},\ldots,\Theta_{n}$ are i.i.d. from an unknown distribution $G$ on $[0,\infty)\times[0,\infty)$. 2. Every player $i$ participates in a random proportion $\lambda_{i}=(\lambda_{i}^{\textnormal{min}},\lambda_{i}^{\textnormal{maj}})$ of all games per minor- and major-league season, and $\lambda_{1},\ldots,\lambda_{n}$ are i.i.d. from an unknown distribution on $[0,\infty)\times[0,\infty)$, and $\lambda_{1},\ldots,\lambda_{n}$ are independent of $\Theta_{1},\ldots,\Theta_{n}$.444 The assumption of independence between $\Theta_{1},\ldots,\Theta_{n}$ and $\lambda_{1},\ldots,\lambda_{n}$ is debatable and warrants further investigation in future work. On the one hand, good batters are more likely to be placed in a batting lineup than bad batters. On the other hand, players who are desired for non-batting skills (catching, fielding, etc.) must be placed in the batting lineup in order to exercise these other skills. 3. For each plater $i$, his RBI count $Z_{i}=(Z_{i}^{\textnormal{min}},Z_{i}^{\textnormal{maj}})$ has conditional distribution $\textnormal{Poi}(\lambda_{i}^{\textnormal{min}}\Theta_{i}^{\textnormal{min}})\otimes\textnormal{Poi}(\lambda_{i}^{\textnormal{maj}}\Theta_{i}^{\textnormal{maj}})$ given $\Theta_{1},\ldots,\Theta_{n},\lambda_{1},\ldots,\lambda_{n}$, and $Z_{1},\ldots,Z_{n}$ are conditionally independent given $\Theta_{1},\ldots,\Theta_{n},\lambda_{1},\ldots,\lambda_{n}$.
Next we describe our denoising methods. All of our methods are based on $G$-modeling (without smoothness assumptions) where we posit that $G$ a bivariate heterogeneous Poisson mixture model with conditionally independent components; this is similar to existing studies on the univariate Poisson case (Polyanskiy and Wu, 2021; Jana et al., 2024) and the multivariate heteroskedastic Gaussian case (Soloff et al., 2024). First, we compute the NPMLE for $G$, yielding an unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$. Second, as outlined in Section D, we compute the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal general-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{GCB}}$ corresponding to constraints on the first and second moment and on component-wise nonnegativity.
The results are shown in Figure 6.2, which we now explain. First, note that the (unconstrained) EB denoiser $\hat{\delta}_{\mathcal{B}}$ reveals a latent structure with a positive trend between the two coordinates and with most players lying below the diagonal dashed line; this reflects the intuitive facts that major league skill is correlated with minor league skill and that most players have lower major league skill than their minor league skill. Second, we compare the unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$ with the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal distribution-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{DCB}}$; the shrinkage in the former leads one to underestimate variability between players. Lastly, compare the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ with the marginal general-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{GCB}}$, and notice that most of the denoised data set is similar, except that $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ produces some negative estimates while $\hat{\delta}_{\mathcal{M}\mathcal{GCB}}$ does not.
Figure 6.2: Denoising the rookie batters’ dataset of Subsection 6.2. We plot the exposure-standardized observations (first), the unconstrained EB denoised data (second), and the marginal variance-constrained EB denoised data (third), and the marginal general-constrained EB denoised data (fourth) corresponding to variance constraints and component-wise nonnegativity constraints.
7 Discussion
Constrained denoising is by now a classical idea that has found use across applications. The optimal transport framework developed in this paper substantially extends the scope of constrained denoising by showing that it is fully compatible with nonparametric empirical Bayes estimators, extending it to higher-dimensional settings, and introducing constraints beyond variance—including distributional and general constraints, as well as conditional constraints under heterogeneity.
For practitioners, several natural questions arise: Which empirical Bayes denoiser should be used among a suite of parametric and nonparametric options? Which type of constraint is most appropriate—variance, distributional, or general? And if general constraints are chosen, which functionals of the latent distribution should be matched? We do not provide definitive answers to these questions, as the optimal choices depend on both the specific application domain and the goals of downstream analyses. Lee et al. (2024) conducted a detailed study providing guidelines for choosing among denoising methods (including variance-constrained approaches) for the analysis of multisite trials. With the expanded framework developed here, we envisage similar guidelines emerging for other concrete application areas, guided by the requirements of domain experts.
From the point of view of statistical theory, an important outstanding question (which is beyond the scope of this paper) concerns lower bounds for our rates of convergence in Section 4. That is, are the rates of convergence in our Theorem 4.3 and Theorem 4.4 optimal? We believe such questions will be interesting for future work, especially in the setting of a nonparametric latent variable distribution and a Gaussian likelihood.
Reproducibility
All figures in the paper can be reproduced in Python with the Jupyter notebooks available at https://github.com/aqjaffe/constrained-denoising-EB-OT.
Acknowledgments
We would like to thank Stephen Raudenbush and Tom Loredo for helpful discussions and for pointing us to the literature on constrained EB estimation. We also thank Brad Ross for providing detailed comments on an early version of this manuscript. Finally, we thank Ellen Woods for many useful conversations about baseball.
Appendix A Additional Empirical Bayes Approaches
While the main body of the paper is primarily focused on nonparametric EB methods, we note that our results apply to (unconstrained) EB denoisers $\hat{\delta}_{\mathcal{B}}$ arising from many different approaches. In this section we briefly review some alternative approaches which we will believe will be important in other applications.
A.1 $F$-Modeling
In some settings, the Bayes denoiser $\delta_{\mathcal{B}}$ depends only on the marginal distribution $F$ of $Z_{1},\ldots,Z_{n}$. Then an EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$ arises by modeling $F$ directly, rather than $G$. This is possible when $P$ is a continuous exponential family with natural parameterization, due to Eddington’s/Tweedie’s formula [Dyson, 1926, Efron, 2011], but it is also possible in many special cases where similar formulas are available [Robbins, 1956, 1982, Cressie, 1982]. Here we follow Efron [2014] and distinguish between $G$-modeling and $F$-modeling approaches, although we note that $F$-modeling can also be viewed as a special case of $G$-modeling.
A.2 Conjugate Parametric Models
Suppose in $d=m=1$ that $\{P_{\theta}\}_{\theta}$ is a mean-parameterized exponential family and that $G$ is a conjugate prior for $\{P_{\theta}\}_{\theta}$. We also assume that $\{P_{\theta}\}_{\theta}$ has a quadratic variance function, meaning there exists a quadratic polynomial $V:\mathbb{R}\to\mathbb{R}$ such that $P_{\theta}$ has variance $V(\theta)$ for all $\theta$; see Morris [1982] for more on such models. It is well-known (see Diaconis and Ylvisaker [1979]) that the Bayes denoiser $\delta_{\mathcal{B}}$ is a linear function in this setting; we have
$\delta_{\mathcal{B}}(z)=\frac{1}{1+\frac{1}{2}V^{\prime\prime}(0)+a^{-2}V(\mu)}z+\left(1-\frac{1}{1+\frac{1}{2}V^{\prime\prime}(0)+a^{-2}V(\mu)}\right)\mu$
where $\mu=\mathbb{E}[\Theta]$ and $a^{2}=\textnormal{Var}(\Theta)$ are the mean and variance of $G$. Although $\mu$ and $a^{2}$ are not known, one can estimate them via the marginal distribution of $Z$ alone, leading to an EB denoiser of the form
$\hat{\delta}_{\mathcal{B}}(z)=\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}z+\left(1-\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}\right)\hat{\mu},$
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{s}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$. One can easily show in this setting that the convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$ has the parametric rate $n^{-1}$.
In this setting, we may use Theorem 3.3 to see that the variance-constrained Bayes denoiser is given explicitly by
$\delta_{\mathcal{VCB}}(z)$ $=\frac{1}{\sqrt{1+\frac{1}{2}V^{\prime\prime}(0)+a^{-2}V(\mu)}}z+\left(1-\frac{1}{\sqrt{1+\frac{1}{2}V^{\prime\prime}(0)+a^{-2}V(\mu)}}\right)\mu$
$=\sqrt{\frac{1-s^{-2}V(\mu)}{1+\frac{1}{2}V^{\prime\prime}(0)}}z+\left(1-\sqrt{\frac{1-s^{-2}V(\mu)}{1+\frac{1}{2}V^{\prime\prime}(0)}}\right)\mu,$
and also that $\hat{\delta}_{\mathcal{VCB}}$ is given by
$\hat{\delta}_{\mathcal{VCB}}(z)=\sqrt{\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}}z+\left(1-\sqrt{\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}}\right)\hat{\mu}.$
In the remainder of this section, we give some explicit formulas for the denoisers above, for the usual parametric models of interest.
Gaussian
As we already discussed in the introduction, a fundamental example is $G=\mathcal{N}(\mu,a^{2})$ and $P_{\theta}=\mathcal{N}(\theta,\sigma^{2})$, where $\mu\in\mathbb{R}$ and $a^{2}>0$ are unknown and $\sigma^{2}>0$ is known. Of course, this fits into the framework above with constant variance function $V\equiv\sigma^{2}.$ As such, the Bayes denoiser is
$\delta_{\mathcal{VCB}}(z)=\frac{a}{\sqrt{a^{2}+\sigma^{2}}}z+\left(1-\frac{a}{\sqrt{a^{2}+\sigma^{2}}}\right)\mu,$
and its EB approximation is
$\hat{\delta}_{\mathcal{VCB}}(z)=\frac{\sqrt{(\hat{s}^{2}-\sigma^{2})_{+}}}{\hat{s}}z+\left(1-\frac{\sqrt{(\hat{s}^{2}-\sigma^{2})_{+}}}{\hat{s}}\right)\hat{\mu}.$
This coincides with the formulas given in Louis [1984] and Ghosh [1992].
Poisson
Suppose that $G$ is a Gamma distribution with unknown shape $k>0$ and scale $\alpha>0$, and that $P_{\theta}=\textnormal{Poi}(\theta)$; this fits into the setting above with linear variance function $V(\theta)=\theta$. Consequently, we can compute the oracle variance-constrained Bayes denoiser and the empirical variance-constrained Bayes denoiser to be
$\delta_{\mathcal{VCB}}(z)=\sqrt{\frac{\alpha}{\alpha+1}}z+\left(1-\sqrt{\frac{\alpha}{\alpha+1}}\right)k\alpha,$
and
$\hat{\delta}_{\mathcal{VCB}}(z)=\frac{\sqrt{(\hat{s}^{2}-\hat{\mu})_{+}}}{\hat{s}}z+\left(1-\frac{\sqrt{(\hat{s}^{2}-\hat{\mu})_{+}}}{\hat{s}}\right)\hat{\mu}.$
Exponential
Suppose $G$ is an inverse Gamma distribution with unknown shape $k>0$ and scale $\alpha>1$, and that $P_{\theta}$ is an exponential distribution with mean $\theta$. (Note that we are using the scale-parameterization of the exponential distribution, rather than the usual rate parameterization.) This fits into the setting above with $V(\theta)=\theta^{2}$. Thus, the oracle and empirical variance-constrained Bayes denoisers are given by
$\delta_{\mathcal{VCB}}(z)=\frac{1}{\sqrt{\alpha}}z+\left(1-\frac{1}{\sqrt{\alpha}}\right)\frac{k}{\alpha-1}$
and
$\hat{\delta}_{\mathcal{VCB}}(z)=\frac{\sqrt{\hat{s}^{2}-\hat{\mu}^{2}}}{\hat{s}\sqrt{2}}z+\left(1-\frac{\sqrt{\hat{s}^{2}-\hat{\mu}^{2}}}{\hat{s}\sqrt{2}}\right)\hat{\mu}.$
respectively.
Geometric
Suppose that $P_{\theta}$ is a Geometric distribution on $\{1,2,\ldots\}$ with success probability equal to $1/(1+\theta)$ so that its variance function is $V(\theta)=\theta+\theta^{2}$. (Note that this is the mean parameterization rather than the natural parameterization.) It is easy to check in this case that the conjugate prior is $\{G_{k,\alpha}\}_{k,\alpha}$ given by
$\frac{\textnormal{d}G_{k,\alpha}}{\textnormal{d}\lambda}(\theta)=\frac{\Gamma(k+\alpha)}{\Gamma(k+1)\Gamma(\alpha-1)}\cdot\frac{\theta^{k}}{(\theta+1)^{k+\alpha}},$
since one can indeed verify that the conditional distribution of $\Theta$ given $\{Z=z\}$ is $G_{k+z-1,\alpha+1}$. We can also compute that, for $\alpha>3$, the first two non-central moments of $G_{k,\alpha}$ are
$\mathbb{E}[\Theta]$ $=\int_{0}^{\infty}\theta\textnormal{d}G_{k,\alpha}(\theta)=\frac{k+1}{\alpha-2}$
$\mathbb{E}[\Theta^{2}]$ $=\int_{0}^{\infty}\theta^{2}\textnormal{d}G_{k,\alpha}(\theta)=\frac{(k+2)(k+1)}{(\alpha-3)(\alpha-2)},$
hence we have
$\textnormal{Var}(\Theta)=\frac{(k+1)(k+\alpha-1)}{(\alpha-3)(\alpha-2)^{2}}.$
In particular, we compute the oracle variance-constrained Bayes denoiser to be
$\delta_{\mathcal{VCB}}(z)=\frac{1}{\sqrt{\alpha-1}}z+\left(1-\frac{1}{\sqrt{\alpha-1}}\right)\frac{k+1}{\alpha-2},$
and the empirical variance-constrained Bayes denoiser to be
$\hat{\delta}_{\mathcal{VCB}}(z)=\frac{\sqrt{(\hat{s}^{2}-\hat{\mu}-\hat{\mu}^{2})_{+}}}{\hat{s}\sqrt{2}}z+\left(1-\frac{\sqrt{(\hat{s}^{2}-\hat{\mu}-\hat{\mu}^{2})_{+}}}{\hat{s}\sqrt{2}}\right)\hat{\mu}.$
Appendix B Discussion of Rates of Convergence
B.1 Variance-Constrained Denoising
This section contains a detailed discussion of the rates of convergence that Theorem 4.3 implies for the convergence of $\hat{\delta}_{\mathcal{VCB}}$ to $\delta_{\mathcal{VCB}}$, where $\delta_{\mathcal{VCB}}$ is the solution to problem ($\mathcal{VCB}$) defined in (3.1) and $\hat{\delta}_{\mathcal{VCB}}$ is its EB approximation defined in Algorithm 1.
First we consider the setting of Gaussian model and nonparametric prior, as in Example 4.1, where we use $G$-modeling to estimate $\hat{\delta}_{\mathcal{B}}$ via the NPMLE. We apply the rates of convergence in [Soloff et al., 2024, Theorem 9] in two further subcases:
  • If $G$ is sub-exponential, then we may take $q=4$ and $\alpha_{n}=n^{-1}(\log n)^{c_{m}}$, for some constant $c_{m}>0$ depending only on the dimension $m$, hence $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate $n^{-1}(\log n)^{c_{m}}$.
  • If $G$ is sufficiently heavy-tailed that (FM) is satisfied only for $2<q<2+\sqrt{m^{2}+4}-m$, then we take $\alpha_{n}=n^{-q(q+m)^{-1}}(\log n)^{c_{m}}$, for some constant $c_{m}>0$ depending only on the dimension $m$, hence $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate $n^{-2(q-2)/q}$.
We emphasize that the results from Soloff et al. [2024] may be used to derive further rates of convergence for $\hat{\delta}_{\mathcal{VCB}}$ under the Gaussian model, but where $G$ has some other structural properties of interest (e.g., sparse support).
Second, we consider the setting of a Poisson model and nonparametric prior, where we can use either the $G$-modeling approach of the NPMLE (Example 4.1) or an $F$-modeling approach (Appendix A.1) in particular the Robbins estimator Robbins [1956]. By the results of Jana et al. [2024], Polyanskiy and Wu [2021], and Shen and Wu [2022], we may derive the following rates:
  • If $G$ is compactly-supported, then we may take $q=4$ and $\alpha_{n}=n^{-1}(\log n/\log\log n)^{2}$, which shows that $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate of convergence $n^{-1}(\log n/\log\log n)^{2}$.
  • If $G$ is sub-exponential, then we may take $q=4$ and $\alpha_{n}=n^{-1}(\log n)^{3}$, so $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate of convergence $n^{-1}(\log n)^{3}$.
  • If $G$ is sufficiently heavy-tailed that (FM) is satisfied only for $42<q<$ 4, then it is known that the $G$-modeling approach leads to the rate $\alpha_{n}=n^{-(q-2)(q+1)^{-1}}(\log n)^{c}$ for some $c>0$ and that the $F$-modeling approach leads to the rate $\alpha_{n}=n^{-(2q-2)(2q+1)^{-1}}(\log n)^{c^{\prime}}$ for some $c^{\prime}>0$. If $2<q<\sqrt{3}+1\approx 2.7321$ then the rate of covariance estimation dominates, and both the $G$-modeling and $F$-modeling approaches lead to $\hat{\delta}_{\mathcal{VCB}}$ achieving the rate $n^{-2(q-2)/q}$.
Despite the fact that the rates above for $G$-modeling and $F$-modeling agree in many cases, we remark that $G$-modeling is generally preferable in practice.
Last, let us consider the setting of an exponential family with its conjugate prior, as in Appendix A.2. As we already discussed, the denoiser $\hat{\delta}_{\mathcal{VCB}}$ from Algorithm 1 has the following formula:
$\hat{\delta}_{\mathcal{VCB}}(z)=\sqrt{\frac{(1-\hat{s}^{-2}V(\hat{\mu}))_{+}}{1+\frac{1}{2}V^{\prime\prime}(0)}}z+\left(1-\sqrt{\frac{(1-\hat{s}^{-2}V(\hat{\mu}))_{+}}{1+\frac{1}{2}V^{\prime\prime}(0)}}\right)\hat{\mu},$
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{\sigma}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$. Since we can take $\alpha_{n}=n^{-1}$ and $q=4$, we obtain the expected behavior, that $\hat{\delta}_{\mathcal{VCB}}$ achieves the parametric convergence rate $n^{-1}$.
B.2 Distribution-Constrained Denoising
This section contains a brief discussion of the rates of convergence yielded by Theorem 4.4 for the convergence of $\hat{\delta}_{\mathcal{DCB}}$ to $\delta_{\mathcal{DCB}}$, where $\delta_{\mathcal{DCB}}$ is the unique solution to problem ($\mathcal{DCB}$) guaranteed by Theorem 3.8 and $\hat{\delta}_{\mathcal{DCB}}$ is its EB estimator defined in Algorithm 2.
First, consider the setting of a Gaussian model and nonparametric compactly supported latent variable distribution $G$. We already saw in Appendix B that assumptions (2M), (ZC), and (BC) are satisfied, and that Soloff et al. [2024, Theorem 9] allows us to take $\alpha_{n}=n^{-1}(\log n)^{c_{m}}$ for some constant $c_{m}>0$ depending only on the dimension $m$. We may also use Soloff et al. [2024, Theorem 11] to take $\beta_{n}=(\log n)^{-1}$ which is known to be the minimax optimal rate for Gaussian deconvolution. (See Dedecker and Michel [2013, Theorem 1].) Lastly, we use Chewi et al. [2025, equation (2.21)] to take
$\gamma_{n}:=\begin{cases}n^{-\nicefrac{{1}}{{2}}}&\textnormal{ if }1\leq m\leq 3\\ n^{-\nicefrac{{1}}{{2}}}\log n&\textnormal{ if }m=4\\ n^{-\nicefrac{{2}}{{m}}}&\textnormal{ if }m\geq 5.\end{cases}$ (B.1)
For any dimension $m$, the deconvolution term dominates and we conclude that $\hat{\delta}_{\mathcal{DCB}}$ achieves the slow rate $(\log n)^{-1}$.
Second, consider the setting of an exponential family with its conjugate prior, as in Subsection A.2. As discussed in Appendix B, assumptions (2M), (ZC), and (BC) are satisfied, and we may take $\alpha_{n}=n^{-1}$. Also, it is expected that one can typically take $\beta_{n}=n^{-1}$. Lastly, note that we may also apply [Chewi et al., 2025, equation (2.21)] to determine that $\gamma_{n}$ is given by (B.1). For any dimension $m$, the empirical Wasserstein distance dominates (since it is always at least as slow as $n^{-\nicefrac{{1}}{{2}}}$), and we conclude that $\hat{\delta}_{\mathcal{DCB}}$ achieves the rate $\gamma_{n}$.
Appendix C Results on General-Constrained Denoising
In this section we detail the oracle- and empirical-level versions of problem ($\mathcal{GCB}$), which were mentioned but not precisely studied in the main body of the paper.
C.1 Oracle General-Constrained Denoising
We begin with the oracle version of problem ($\mathcal{GCB}$) of general-constrained Bayes denoising. That is, we fix a linearly-independent collection $\psi_{1},\ldots,\psi_{k}:\mathbb{R}^{m}\to\mathbb{R}$ of continuous functions (which are also assumed to be $G$-integrable). For convenience, we write $\Psi:=\{\psi_{1},\ldots,\psi_{k}\}$. Roughly speaking, any constraint on the distribution of $\delta(Z)$ can be cast in the form above (although we may need to take $k$ to be infinite); indeed, ($\mathcal{GCB}$) reduces to ($\mathcal{B}$) if $\Psi$ is empty, ($\mathcal{VCB}$) if $\Psi$ consists of all quadratic forms, and ($\mathcal{DCB}$) if $\Psi$ consists of all bounded continuous functions. In addition to subsuming these earlier examples, we are motivated by the following.
Example C.1 (Higher moment constraints).
In the univariate case $m=d=1$ and for some fixed positive integer $k$, one may aim to minimize the denoising risk while matching the first $k$ moments of $\delta(Z)$ to those of $\Theta$, thereby solving the problem
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[|\delta(Z)-\Theta|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}\to\mathbb{R}\\ \textnormal{with}&\mathbb{E}\left[(\delta(Z))^{\ell}\right]=\mathbb{E}\left[\Theta^{\ell}\right]\quad\textnormal{ for all }\ell=1,\ldots,k.\end{cases}$ (C.1)
This, of course, coincides with the case of variance constraints when $k=2$, and higher moment constraints force the distribution of $\delta(Z)$ to be even closer to the distribution $G$ of $\Theta$. Along these lines it is shown in Armstrong et al. [2022], in the context of constructing nonparametric EB confidence intervals, that imposing constraints on the fourth moment leads to EB procedures which nearly adapt to the parametric (Gaussian) setting.
Example C.2 (Errors-in-variables regression continued: moment adjusted imputation).
Bay [1997] and Thomas et al. [2011] propose an intermediate approach called moment adjusted imputation that bridges the methods described in Examples 3.2 and 3.7. Their approach uses a denoiser that matches a controlled number of moments and cross-moments between $X$ and $Y$, extending beyond the two moments in moment reconstruction but without enforcing full distributional equivalence. For instance, when $X\in\mathbb{R}$, they aim to estimate $X$ while preserving multiple moments such as $\mathbb{E}[X]$, $\mathbb{E}[X^{2}]$, $\mathbb{E}[X^{3}]$, $\mathbb{E}[X^{4}]$, as well as cross-moments like $\mathbb{E}[XY]$ and $\mathbb{E}[X^{2}Y]$. Such a denoiser is given by problem ($\mathcal{GCB}$) with the choices $Z=(W,Y)$, $\Theta=(X,Y)$, $\psi_{\ell}(x,y)=x^{\ell}$ for $\ell=1,\dotsc,4$, and $\psi_{5}(x,y)=xy$, $\psi_{6}(x,y)=x^{2}y$.
We need some further assumptions and notation in order to state our main result on problem ($\mathcal{GCB}$). First, let us define $\Gamma(F;G,\Psi)$ to be the set of all joint distributions $\pi$ on $\mathbb{R}^{d}\times\mathbb{R}^{m}$ such that if $(Z,D)\sim\pi$ then $Z\sim F$ and $\mathbb{E}[\psi_{\ell}(D)]=\mathbb{E}[\psi_{\ell}(\Theta)]$ for all $1\leq\ell\leq k$.
Furthermore, we assume that
[…]
We note that (C) is a mild condition, and that it is satisfied whenever $\{P_{\theta}\}_{\theta}$ is a suitable exponential family, regardless of $G$ [García-Trillos and Sen, 2024, Remark 2.2]; for example, it holds for Gaussian or Poisson families. (It holds trivially whenever $F$ has discrete support.) Lastly, we assume that
[…]
Here, a function $\psi:\mathbb{R}^{m}\to\mathbb{R}$ is said to have at least quadratic growth if there exist constants $C,R>0$ such that we have $\|\eta\|^{2}\leq C\psi(\eta)$ for all $\eta\in\mathbb{R}^{m}$ with $\|\eta\|>R$.
Finally, we get the following. We remind the reader that $\delta_{\mathcal{B},H}$ is the Bayes denoiser when $\Theta$ comes from distribution $H\in\mathcal{P}(\mathbb{R}^{m})$, and that the cost function $c_{G}(\cdot,\cdot)$ is defined in (3.3).
Theorem C.3.
Under assumptions (ZC), (BC), (C), and (QG), the problem
$\underset{\pi\in\Gamma(F;G,\Psi)}{\textnormal{minimize}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\,\textnormal{d}\pi(z,\eta),$ (C.2)
has a unique solution $\pi_{\mathcal{GCB}}$, and this solution is concentrated on the graph of a function which we denote $\delta_{\mathcal{GCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. If $H$ denotes the marginal distribution of $D$ when $(Z,D)\sim\pi_{\mathcal{GCB}}$, then $\delta_{\mathcal{GCB}}$ can be written as $\delta_{\mathcal{GCB}}=\nabla\phi\circ\delta_{\mathcal{B},H}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, and we have $R_{\mathcal{GCB}}=R_{\mathcal{B}}+W_{2}^{2}(H,(\delta_{\mathcal{B}})_{\#}F)$. Consequently, (C.2) is a tight convex relaxation of ($\mathcal{GCB}$), and $\delta_{\mathcal{GCB}}$ is the $F$-a.s. unique solution to ($\mathcal{GCB}$).
As in the case of problem ($\mathcal{DCB}$), this result provides conditions under which problem ($\mathcal{GCB}$) has a unique solution which can be found by solving the linear program (C.2). We refer to the probability measure as $H$ as the projected latent variable distribution since, among all probability measures whose integrals with $\Psi$ match those of $G$, it is the nearest to $(\delta_{\mathcal{B}})_{\#}F$ with respect to the distance $W_{2}$.
Example C.4 (Higher moment constraints continued).
To provide further intuition on the projected latent variable distribution $H$, we characterize it in some particular cases of Example C.1. If $k=0$ or $k=1$, then $\delta_{\mathcal{B}}$ is feasible hence optimal for (C.1), so $H$ equals $(\delta_{\mathcal{B}})_{\#}F$. If $k=2$, then (C.1) is equivalent to ($\mathcal{VCB}$), so $H$ equals the pushforward of $F$ by an affine function of $\delta_{\mathcal{B}}$. If $G$ is compactly supported and $H_{k}$ denotes the projected latent variable distribution for constraints on $k$ moments, then $H_{k}$ converges weakly to $G$ as $k\to\infty$.
C.2 Empirical General-Constrained Denoising
Since in Theorem C.3 we provided sufficient conditions for the existence of a unique solution to problem ($\mathcal{GCB}$), we presently develop an EB methodology for targeting this oracle denoiser. Methodologically, we regard ($\mathcal{GCB}$) as a way to tradeoff between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$); the former only encodes limited information about the latent variable distribution (the first two moments), the latter suffers from slow rates of convergence (due to the difficulty of the deconvolution problem) and may have larger denoising risk. Often some intermediate behavior may be desired.
We describe our proposed procedure in Algorithm 3, which is similar to the approach we previously used for the distribution-constrained problem ($\mathcal{DCB}$). Since we can estimate the marginal distribution $F$ via $\bar{F}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{Z_{i}}$ and the cost function $c_{G}$ via $\hat{c}_{n}$ through $\hat{\delta}_{\mathcal{B}}$, it only remains to approximate the constraints of the linear program (C.2). Assuming that this is possible (that is, that the integrals of $\Psi$ with $\hat{G}_{n}$ converge to the integrals of $\Psi$ with $G$), we may let $\hat{\delta}_{\mathcal{GCB}}$ denote the barycentric projection of an optimal solution to the corresponding linear program. Note also that we restrict the $\eta$ variable of the linear program to lie in a compact set $K\subseteq\mathbb{R}^{m}$ (line 6).
In order to give a rate of convergence of $\hat{\delta}_{\mathcal{GCB}}$ to $\delta_{\mathcal{GCB}}$, we need to present a set of assumptions. To set this up, we first assume (ZC), (BC), (C), and (QG) so that Theorem C.3 is in effect; then, we recall that there is a well-defined projected latent variable distribution which we denote $H$. Second, we assume
the support of $\textnormal{the support of }H\textnormal{ is contained in the compact set }K\subseteq\mathbb{R}^{m}$ (CS)
which implies that $H$ is compactly supported and that the support constraint of Algorithm 3 is well-specified. Third, we need
$\begin{pmatrix}\int_{\mathbb{R}^{m}}\psi_{1}\textnormal{d}G\\ \vdots\\ \int_{\mathbb{R}^{m}}\psi_{k}\textnormal{d}G\end{pmatrix}\in\left\{\begin{pmatrix}\int_{\mathbb{R}^{m}}\psi_{1}\textnormal{d}G^{\prime}\\ \vdots\\ \int_{\mathbb{R}^{m}}\psi_{k}\textnormal{d}G^{\prime}\end{pmatrix}:G^{\prime}\in\mathcal{P}(\mathbb{R}^{m})\right\}^{\circ}=:\mathcal{M}^{\circ}(\Psi).$ (INT)
where $A^{\circ}$ denotes the interior of a set $A\subseteq\mathbb{R}^{k}$; this condition states that the integrals of $\psi_{1},\ldots,\psi_{k}$ under the latent variable distribution $G$ do not lie on the boundary of all possible values of the integrals. Fourth, we impose assumption (EC) and that $\hat{G}_{n}$ therein (typically taken to be the NPMLE or a smoothed version thereof) additionally satisfies
$\int_{\mathbb{R}^{m}}\psi_{\ell}(\eta)\textnormal{d}\hat{G}_{n}(\eta)\to\int_{\mathbb{R}^{m}}\psi_{\ell}(\eta)\textnormal{d}G(\eta)\textnormal{ for all }1\leq\ell\leq k$ (CC)
in probability as $n\to\infty$. For $\psi_{\ell}$ with at most quadratic growth, such convergence follows from (EC) because of Villani [2009, Definition 6.8]. For $\psi_{\ell}$ whose growth is polynomial of degree $p\geq 2$, it is sufficient to additionally require $W_{p}(\hat{G}_{n},G)=o_{\mathbb{P}}(1)$ on top of (EC).
Next we introduce some assumptions which are analogous to assumptions used in Theorem 4.4. First, write $\hat{H}_{n}$ for the distribution of the $\mathbb{R}^{m}$-coordinate under the optimal coupling $\hat{\pi}_{\mathcal{GCB}}$ from Algorithm 3. We require that, for some specified sequence $\delta_{n}\to 0$, we have
$W_{2}^{2}(\hat{H}_{n},H)=O_{\mathbb{P}}(\delta_{n}).$ (D)
It will be shown during the course of our main theorem that (D) follows from (CS), (INT), and (CC), so this assumption only amounts to quantifying the rate of convergence. Second, we assume
$\delta_{\mathcal{GCB}}=\nabla\phi\circ\delta_{\mathcal{B},H}\textnormal{ where}\phi\textnormal{ is $\lambda$-strongly convex and $L$-smooth for some }\lambda,L>0,$ (R)
which is analogous to assumption (R) but with $\delta_{\mathcal{B},H}$ in place of $\delta_{\mathcal{B},G}\equiv\delta_{\mathcal{B}}$. Lastly, we make the following assumption on the convergence of the empirical Wasserstein distance for i.i.d. samples $\eta_{1},\ldots,\eta_{n}$ from $H$ defined via $\eta_{i}:=\delta_{\mathcal{GCB}}(Z_{i})$: for some specified sequence $\varepsilon_{n}\to 0$ we have
$W_{2}^{2}\left(\bar{H}_{n},\;H\right)=O_{\mathbb{P}}(\varepsilon_{n})\qquad\textnormal{ where }\qquad\bar{H}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\eta_{i}}$ (EW)
which is analogous to (EW).
Algorithm 3 An EB estimator $\hat{\delta}_{\mathcal{GCB}}$ of the oracle general-constrained Bayes denoiser $\delta_{\mathcal{GCB}}$.
1:procedure GeneralConstrainedEB($Z_{1},\ldots,Z_{n};\{P_{\theta}\}_{\theta}$)
2:     input: samples $Z_{1},\ldots,Z_{n}$, likelihood $\{P_{\theta}\}_{\theta}$, compact $K\subseteq\mathbb{R}^{m}$
3:     output: denoising function $\hat{\delta}_{\mathcal{GCB}}:\{Z_{1},\ldots,Z_{n}\}\to\mathbb{R}^{m}$
4:     $\hat{\delta}_{\mathcal{B}}(\,\cdot\,)\leftarrow$ EB approximation of $\delta_{\mathcal{B}}(\,\cdot\,)$
5:     $\hat{G}_{n}\leftarrow$ estimate of $G$
6:     $\hat{c}_{n}(Z_{i},\eta)\leftarrow\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\eta\|^{2}$ for all $1\leq i\leq n$ and $\eta\in K$
7:     $\hat{\pi}_{\mathcal{GCB}}\leftarrow$ minimize $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}(z,\eta)\textnormal{d}\pi(z,\eta)$
8:                 over         probability measures $\pi\in\mathcal{P}(\mathbb{R}^{d}\times K)$
9:                 with        $\pi(\{Z_{i}\}\times K)=\frac{1}{n}$ for all $1\leq i\leq n$
10:                 and         $\int_{\mathbb{R}^{d}\times K}\psi_{\ell}(\eta)\textnormal{d}\pi(z,\eta)=\int_{K}\psi_{\ell}\textnormal{d}\hat{G}_{n}$ for all $1\leq\ell\leq k$
11:     $\hat{\delta}_{\mathcal{GCB}}(Z_{i})\leftarrow\int_{K}\eta\,\textnormal{d}\hat{\pi}_{\mathcal{GCB}}(\eta\,|\,Z_{i})$ for all $1\leq i\leq n$
12:     return $\hat{\delta}_{\mathcal{GCB}}$
13:end procedure
Finally, these considerations lead us to the following result.
Theorem C.5.
Under assumptions (ZC), (BC), (C), (QG), (CS), (INT), (CC), (D), (R), (EW), and (EBQ), the denoiser $\hat{\delta}_{\mathcal{GCB}}$ from Algorithm 3 satisfies
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{GCB}}(Z_{i})-\delta_{\mathcal{GCB}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\delta_{n}\vee\varepsilon_{n})$ (C.3)
as $n\to\infty$.
The conclusion of Theorem C.5 for the denoiser $\hat{\delta}_{\mathcal{GCB}}$ given in Algorithm 3 is that the rate of convergence of general-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising, convergence of the projected latent variable distributions, and convergence of the empirical Wasserstein distance is slowest. The rate of convergence of the projected latent variable distributions, described by $\delta_{n}$, exactly quantifies the statistical tradeoff of interpolating between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$). In the cases we consider in Example C.4, we may anticipate $\delta_{n}$ to decay at nearly parametric rates when $k=0$ (no constraints), or when $k=2$ (variance constraints) and the empirical Bayes denoiser may be estimated at fast rates (as discussed in Example 4.1), or inversely polylogarithmic in $n$ when $k\to\infty$ (and we effectively need to deconvolve the latent variable distribution).
Appendix D Details for Heterogeneous (Empirical) Constrained Bayes
The unconstrained denoising problem in model (5.1) is as follows:
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z;\Xi)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\times\mathbb{R}^{p}\to\mathbb{R}^{m}.\end{cases}$ (D.1)
Its solution is the conditional mean of $\Theta$ given $(Z,\Xi)$, which we denote by
$\delta_{\mathcal{B}}(z;\xi)=\mathbb{E}[\Theta\,|Z=z,\Xi=\xi].$
In the next subsections, we consider adding constraints to problem (D.1). While we will not prove rigorous results regarding asymptotic theory, we extend the strategies of the previous sections to accommodate the heterogeneity. An important feature of heterogeneous constrained denoising is that we now must distinguish between two kinds of constraints: the so-called marginal constraints on the distribution of $\delta_{\mathcal{B}}(Z;\Xi)$, and the so-called conditional constraints on the distribution of $\delta_{\mathcal{B}}(Z;\Xi)$ given $\{\Xi=\xi\}$.
Before explaining how to extend our constrained denoising methods to model (5.1), we first discuss possible applications of the model in more detail.
D.1 Applications of the heterogeneous setting
Model (5.1), allows for heterogeneity in both the prior $(G_{\xi})$ and the likelihood $P_{\theta,\xi}$. Two important examples for heterogeneity in the likelihood are the following:
  • (Heteroskedastic Gaussian.) For $m=d$, assume that the parameter $\theta$ is a vector in $\mathbb{R}^{m}$ and the parameter $\xi$ is a positive semi-definite $m\times m$ matrix, which we choose to denote by $\Sigma$. Let $P_{\theta,\Sigma}=\mathcal{N}(\theta,\Sigma)$. As one concrete application, suppose that conditional on $\Theta_{i}$, $Z_{i}$ is a sample average of $K_{i}$ i.i.d. observations $Z_{ij}$ with $Z_{ij}\sim\mathcal{N}(\Theta_{i},\sigma^{2})$. Then $Z_{i}\sim\mathcal{N}(\Theta_{i},\sigma_{i}^{2})$ with $\sigma_{i}^{2}=\sigma^{2}/K_{i}$. The heteroskedastic Gaussian setting encompasses this situation by taking $m=d=1$ and $\Xi_{i}=\sigma^{2}/K_{i}$.
  • (Poisson with heterogeneous exposure.) For $m=d=1$, the parameter $\theta$ is a non-negative real number and the parameter $\xi$ is a non-negative real number, which we choose to denote by $\lambda$. Then let $P_{\theta,\lambda}=\textnormal{Poi}(\lambda\theta)$. We call $\lambda$ the exposure [Norberg, 1979]. The intuition here is that we are interested in the mean of the $i$-th observation normalized by the exposure time, that is, $\Theta_{i}=\mathbb{E}[Z_{i}/\lambda_{i}]$; see Subsection 6.2 for an application to baseball in which $\lambda_{i}$ is the number of games played by the $i$-th player.
Moreover, in several applications, the distribution of $\Theta$ may depend on $\Xi$. Such dependence implies that $\Theta_{1},\ldots,\Theta_{n}$ are no longer exchangeable. Examples where this situation occurs are as follows.
  • (Heteroskedastic Gaussian, continued.) In our application to astronomy (Subsection 6.1), we assume that $\Theta$ is independent of $\Xi\equiv\Sigma$, and as mentioned therein, this is a common assumption in the literature. However, several authors have noted that this assumption can be violated [Stephens, 2017, Gu and Koenker, 2017b, Weinstein et al., 2018, Chen, 2022, Ignatiadis and Sen, 2025]. If we model $G_{\xi}$ as a function of $\xi$ (using, e.g., the G-modeling approach of Chen [2022]), then we can address this issue (and the general model in (5.1) accommodates this dependence).
  • (Side-information.) In several common applications, we observe covariates (features) $X_{i}$ in addition to the noisy observations $Z_{i}$ for each $i$. The covariates $X_{i}$ are informative about $\Theta_{i}$, though commonly they do not directly impact the likelihood, that is, $Z_{i}$ is independent of $X_{i}$ conditionally on $\Theta_{i}$. The setting with side-information falls under our framework by letting $\Xi_{i}=X_{i}$ and $P_{\theta,\xi}=P_{\theta}$. Examples of empirical Bayes modeling with side-information include Fay III and Herriot [1979], Cohen et al. [2013], Ignatiadis and Wager [2019], Banerjee et al. [2023].
  • (Heteroskedasticity and side-information.) It is possible to have both side-information and heterogeneity in the likelihood at the same time, such as described in Chen [2022], Kwon [2023], Ghosh et al. [2025]. We can model this situation, e.g., in the heteroskedastic Gaussian setting, by letting $\Xi_{i}=(\Sigma_{i},X_{i})$. Herein, $P_{\theta,\xi}$ only depends on the likelihood covariance (not on the side-information), while the prior can depend on both $\Sigma_{i}$ and $X_{i}$.
D.2 Variance-Constrained Denoising
We first seek to extend the variance-constrained denoising problem ($\mathcal{VCB}$) to the heterogeneous setting and consider two approaches; imposing a marginal variance constraint and imposing a conditional variance constraint.
D.2.1 Marginal Variance-Constrained Denoising
For the first approach, we require that the marginal mean and variance of the denoised data match the marginal mean and variance of the latent variable. That is, we solve problem (D.1) subject to the constraints
$\mathbb{E}[\delta(Z;\Xi)]=\mathbb{E}[\Theta]\qquad\textnormal{ and }\qquad\textnormal{Cov}(\delta(Z;\Xi))=\textnormal{Cov}(\Theta).$ (D.2)
This approach was also followed by Ghosh [1992] in a small area application with heterogeneous sample sizes across areas. By noticing that this is closely related to the setting of Section 3.1 but with observation $\tilde{Z}:=(Z,\Xi)$, it can be shown that the oracle denoiser is
$\delta_{\mathcal{M}\mathcal{VCB}}(z;\xi):=\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z;\Xi))}^{\textnormal{Cov}(\Theta)}(\delta_{\mathcal{B}}(z;\xi)-\mu)+\mu$
where $\mu=\mathbb{E}[\Theta]$. In order to implement this in practice, we can apply Algorithm 1 but with the requisite terms $\mathbb{E}[\Theta]$, $\textnormal{Cov}(\Theta)$, and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z;\Xi))$ estimated through modified approaches. It is clear that the third term can be estimated via the sample covariance of $\hat{\delta}_{\mathcal{B}}(Z_{1};\Xi_{1}),\ldots,\hat{\delta}_{\mathcal{B}}(Z_{n};\Xi_{n})$ where $\hat{\delta}_{\mathcal{B}}(\,\cdot\,;\,\cdot\,)$ is any EB denoiser which accommodates the heterogeneity. When $\hat{\delta}_{\mathcal{B}}(\,\cdot\,;\,\cdot\,)$ arises from $G$-modeling, then one has an estimate $\hat{G}_{n}$ of $G$ and one can estimate $\mathbb{E}[\Theta]$ and $\textnormal{Cov}(\Theta)$ via
$\hat{\mu}:=\int_{\mathbb{R}^{m}}\theta\textnormal{d}\hat{G}_{n}(\theta)\qquad\text{ and}\qquad\,\hat{A}:=\int_{\mathbb{R}^{m}}(\theta-\hat{\mu})(\theta-\hat{\mu})^{\top}\textnormal{d}\hat{G}_{n}(\theta).$ (D.3)
D.2.2 Conditional Variance-Constrained Denoising
Another possible constraint is that the conditional mean and covariance of $\delta(Z;\Xi)$ match the conditional mean and covariance of $\Theta$ given $\Xi$. In other words, we may solve problem (D.1) subject to the constraints
$\mathbb{E}[\delta(Z;\Xi)\,|\,\Xi]=\mathbb{E}[\Theta\,|\,\Xi]\qquad\textnormal{and }\qquad\textnormal{Cov}(\delta(Z;\Xi)\,|\,\Xi)=\textnormal{Cov}(\Theta\,|\,\Xi)\qquad\textnormal{ almost surely.}$. (D.4)
In this setting, it can be shown that the oracle denoiser is
$\delta_{\mathcal{C}\mathcal{VCB}}(z,\xi)=\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z;\Xi)\,|\,\Xi=\xi)}^{\textnormal{Cov}(\Theta\,|\,\Xi=\xi)}(\delta_{\mathcal{B}}(z;\xi)-\mu)+\mu,$
where $\mu=\mathbb{E}[\Theta]$. Our EB approach will target the denoiser $\delta_{\mathcal{C}\mathcal{VCB}}$ by plugging in suitable estimates for $\mathbb{E}[\Theta],\textnormal{Cov}(\Theta\,|\,\Xi=\xi)$, and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z,\Xi)\,|\,\Xi=\xi)$. If we additionally assume $\Theta\perp\!\!\!\!\perp\Xi$ (as we will assume in our later applications), then we may estimate $\mathbb{E}[\Theta]$ and $\textnormal{Cov}(\Theta)\equiv\textnormal{Cov}(\Theta\,|\,\Xi)$ in the same way as above, via the corresponding moments of the NPMLE given in (D.3). However, $\textnormal{Cov}(\delta_{\mathcal{B}}(Z,\Xi)\,|\,\Xi=\xi)$ must be handled more carefully. If we are in the setting of $G$-modeling and the measure $\hat{G}_{n}$ is finitely supported (which is true for the NPMLE), then we may approximate
$\textnormal{Cov}(\delta_{\mathcal{B}}(Z;\Xi)\,|\,\Xi=\xi)=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{d}}(\delta_{\mathcal{B}}(z;\xi)-\mu)(\delta_{\mathcal{B}}(z;\xi)-\mu)^{\top}P_{\theta,\xi}(\textnormal{d}z)\textnormal{d}G(\theta).$
via
$\sum_{j=1}^{r}w_{j}\int_{\mathbb{R}^{d}}(\hat{\delta}_{\mathcal{B}}(z,\xi)-\mu)(\hat{\delta}_{\mathcal{B}}(z,\xi)-\mu)^{\top}P_{\theta_{j},\xi}(\textnormal{d}z).$ (D.5)
whenever $\hat{G}_{n}=\sum_{j=1}^{r}w_{j}\delta_{\theta_{j}}$. Lastly, the finitely-many integrals appearing on the right side may be estimated in many different ways, since the likelihood $\{P_{\theta,\xi}\}_{\theta}$ is known, e.g., by numerical integration, or by Monte Carlo.
D.2.3 Numerical Illustration
In order to highlight the difference between the marginal variance-constrained and conditional variance-constrained denoisers $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}$, we consider a simulation in the univariate heteroskedastic Gaussian setting.
More precisely, we suppose that $\sigma^{2}_{1},\ldots,\sigma^{2}_{n}$ are i.i.d. which are equal to $0.5$ or $8$ with equal probability, and we let $\Theta_{1},\ldots,\Theta_{n}$ be i.i.d from a standard Gaussian distribution $\mathcal{N}(0,1)$. Conditional on these values, we let $Z_{1},\ldots,Z_{n}$ be independent with $Z_{i}$ possessing distribution $\mathcal{N}(\theta_{i},\sigma_{i}^{2})$ for all $1\leq i\leq n$ with $n=1,500$. An observation $Z_{i}$ is called a low-variance observation if $\sigma^{2}_{i}=0.5$ and it is called a high-variance observation if $\sigma^{2}_{i}=8$.
In Figure D.1 we show the result of applying three denoisers to this data set. The marginal variance-constrained denoiser (third column) leads to the incorrect variance when applied only to the low-variance observations (second row) or the high-variance observations (third row). Contrarily, the conditional variance-constrained denoiser (fourth column) obtains the correct variance for both subsets of the data.
D.3 Marginal Distribution-Constrained Denoising
Next we seek to extend the distribution-constrained problem ($\mathcal{DCB}$) to the heterogeneous setting. To state the problem precisely, we aim to solve problem (D.1) subject to the constraint that the marginal distribution of the denoised data matches the marginal distribution of the latent variable, meaning
$\delta(Z;\Xi)\overset{\mathcal{D}}{=}\Theta.$ (D.6)
As before, we observe that this is similar to Section 3.2 with observation $\tilde{Z}:=(Z,\Xi)$, hence the oracle denoiser is the composition of the Bayes denoiser $\delta_{\mathcal{B}}$ with the OT map from the distribution of $\delta_{\mathcal{B}}(Z,\Xi)$ to the distribution $G$ of $\Theta$. In order to modify Algorithm 2 to implement this, we need to replace $Z_{i}$ with $(Z_{i},\Xi_{i})$ everywhere. For example, we replace the definition of the cost function (line 7) by
$\hat{c}_{n}((Z_{i},\Xi_{i}),\eta)\leftarrow\|\hat{\delta}_{\mathcal{B}}(Z_{i},\Xi_{i})-\eta\|^{2}\quad\textnormal{ for all }1\leq i\leq n\quad\textnormal{ and }\eta\in\mathbb{R}^{m}$
and we replace barycentric projection (line 12) by
$\hat{\delta}_{\mathcal{M}\mathcal{DCB}}(Z_{i},\Xi_{i})\leftarrow\int_{\mathbb{R}^{m}}\eta\,\textnormal{d}\hat{\pi}_{\mathcal{M}\mathcal{DCB}}(\eta\,|\,(Z_{i},\Xi_{i}))\quad\textnormal{ for all }1\leq i\leq n,$
where $\hat{\pi}_{\mathcal{M}\mathcal{DCB}}$ is the solution to the resulting OT problem, which is still a linear program.
While we focus only on the case of marginal distribution constraints, we remark that an extension to conditional distributional constraints is likely possible, but with further statistical and computational complications.
Figure D.1: Denoising in the heterogeneous setting of Section D to a toy data set. By rows, we show histograms of all of the observations (first row), the subset of low-variance observations (second row), and the subset of high-variance observations (third row). By columns, we show the raw data (first column), the EB denoised data set (second column), the marginal variance-constrained EB denoised data set (third column), and the conditional variance-constrained EB denoised data set (fourth column). In each plot we show the density of the latent variable distribution (blue) for convenience.
D.4 Marginal General-Constrained Denoising
Lastly, we consider general constraints on the marginal distribution of $\delta(Z,\Xi)$, meaning we require
$\mathbb{E}[\psi_{\ell}(\delta(Z,\Xi))]=\mathbb{E}[\psi_{\ell}(\Theta)]\quad\textnormal{ for all }1\leq\ell\leq k,$ (D.7)
where $\psi_{1},\ldots,\psi_{k}$ are some fixed functions. As in Subsection 3.3, the oracle denoiser in this problem is the composition of the Bayes denoiser $\delta_{\mathcal{B}}$ with an OT map from the distribution of $\delta_{\mathcal{B}}(Z,\Xi)$ to some distribution $H$ (whose integrals with respect to $\Psi$ agree with those of $G$) but which is not a priori known. Nonetheless, this denoiser can be approximated via applying the same modifications of Subsection D.3 to Algorithm 3.
Appendix E Computational Considerations
In this section, we briefly describe some computational considerations that arise in the implementation of various algorithms described in the main body of the paper. Our procedures also inherit some computational complexity from the unconstrained EB denoisers which they take as input. Standard OT problems are solved with the Python Optimal Transport (POT) package [Flamary et al., 2021], and other convex optimization problems are solved via CVXPY [Diamond and Boyd, 2016] using the Mosek  [ApS, 2025] interior point convex programming solver.
E.1 $G$-modeling via nonparametric maximum likelihood
Our $G$-modeling methods with a discrete NPMLE are based on the software implementation in the Python package npeb [^5] Available at: https://github.com/jake-soloff/npeb. This implementation approximates the NPMLE through a two-step procedure:
  1. 1.
    Computing the NPMLE over a fixed discrete set of atoms using CVXPY/Mosek following the framework of Koenker and Mizera [2014] (cf. Soloff et al. [2024] for a theoretical study of discretization considerations).
  2. 2.
    Running expectation-maximization to improve atoms and weights jointly.
In the Gaussian models (Figure 1.1, Figure D.1, and Figure 6.1) we inialize the NPMLE atoms in Step 1 to be the data points themselves, following the so-called exemplar$+$ method of Soloff et al. [2024]. In the Poisson model of Figure 6.2 we initialize the NPMLE atoms of step 1 to be a $50\times 50$ equispaced grid for the minimal bounding box of the data.
For our $G$-modeling (e.g., in Section 6.1) with a smooth prior $G$ that is a Gaussian mixture model with a lower-bounded covariance matrix, we employ the strategy of Magder and Zeger [1996, Section 2]. This approach reduces the computation of the NPMLE over the mixture class to that of the discrete NPMLE, allowing us to again apply the npeb package. Here we also choose the number of grid points and their location for Step 1. of npeb to be a $150\times 150$ equispaced grid for the minimal bounding box of the data.
E.2 Variance-Constrained Denoising
First we consider problem ($\mathcal{VCB}$) on denoising subject to variance constraints. Because of the explicit formulas afforded by the geometry of the Bures-Wasserstein space, we can avoid numerically solving an optimization problem in this case. Indeed, each step of Algorithm 1 has an explicit formulation (except for line 4, which can be implemented by existing EB approaches).
Second, we consider the two variance-constrained denoising problems in the heterogeneous setting of Section D. The marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ does not require any further computational burden compared to $\hat{\delta}_{\mathcal{VCB}}$, at least in the case where $\hat{\delta}_{\mathcal{B}}$ arises from $G$-modeling. This is the case in all of our applications except for Figure D.1. The conditional variance-constrained denoiser $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}$ is more complicated since it requires the numerical integration of the terms in (D.5) for each evaluation of $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}(Z_{i},\Xi_{i})$; we compute these integrals via Monte Carlo with 100 trials each.
E.3 Distribution- and General-Constrained Denoising
Our procedures Algorithm 2 and Algorithm 3 for approximating the solution to the distribution-constrained problem ($\mathcal{DCB}$) and the general-constrained problem ($\mathcal{GCB}$) both require numerically solving a linear program (lines 7 to 10), so we briefly discuss some details of this. For problem ($\mathcal{DCB}$) we use the solver from POT which is specifically designed to handle OT problems, and for problem ($\mathcal{GCB}$) we use Mosek’s general-purpose solvers for linear programming in CVXPY. In either case, some discretization of the $\eta\in\mathbb{R}^{m}$ coordinate is necessary. We always choose the discretization to be an equispaced grid on some axis-aligned bounding box, and the main difficulty is choosing an appropriate discretization.
In the cases where $m=1$ (i.e., Figure D.1), the complexity of the linear program problem scales linearly in the number of discretization points $k$. So, it is possible to take $k$ quite large without computational burden; for the sake of simplicity. In these cases, we fixed the discretization of $\eta$ to consist of $k=200$ equispaced points in the minimal bounding interval of the observations.
In the cases where $m\geq 2$ (i.e., Figure 1.1, Figure 6.2, and Figure 6.1) the computation is more burdensome. We take $k=300$ in problem ($\mathcal{DCB}$) since the POT package is highly optimized for this task, and we take $k=200$ for problem ($\mathcal{GCB}$) although computation is a bit slow.
Additionally, we remark that naively discretizing the linear program in Algorithm 3 may lead to infeasibility; this is because $\hat{G}_{n}$ witnesses feasibility for arbitrary probability measures but not necessarily for probability measures supported on the prescribed grid points. To avoid this, we augment the discretization by adding the NPMLE atoms in addition to the grid points.
Appendix F Proofs from Section 3
This section contains the proofs of the main results in Section 3 on the population-level denoising problems.
Proof of Theorem 3.3.
We begin by performing some manipulation on the objective of ($\mathcal{VCB}$). Note that by assumption (2M) that all terms below are well-defined. For feasible $\delta$, let us expand
$\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]=\mathbb{E}\left[\left\|\delta_{\mathcal{B}}(Z)-\Theta+\delta(Z)-\delta_{\mathcal{B}}(Z)\right\|^{2}\right]$
$=\mathbb{E}\Big{[}\left\|\delta_{\mathcal{B}}(Z)-\Theta\right\|^{2}\Big{]}+\mathbb{E}\Big{[}\left\|\delta(Z)-\delta_{\mathcal{B}}(Z)\right\|^{2}\Big{]}+2\mathbb{E}\Big{[}\left(\delta_{\mathcal{B}}(Z)-\Theta\right)^{\top}(\delta(Z)-\delta_{\mathcal{B}}(Z))\Big{]}.$
Notice that the first term on the right side is exactly the Bayes risk and that it is irreducible since it does not depend on $\delta$. The second term on the right side is exactly equal to
$\mathbb{E}\left[\left\|\delta(Z)-\delta_{\mathcal{B}}(Z)\right\|^{2}\right]=\textnormal{Var}\left(\delta(Z)-\delta_{\mathcal{B}}(Z)\right),$
where
$\textnormal{Var}(X):=\textnormal{tr}(\textnormal{Cov}(X))=\sum_{i=1}^{m}\textnormal{Var}(X_{i})$
denotes the variance of a random vector $X=(X_{1},\ldots,X_{m})$, since $\delta$ being feasible implies $\mathbb{E}[\delta(Z)-\delta_{\mathcal{B}}(Z)]=0$. Finally, the third term on the right vanishes, since we can use the law of total expectation to compute:
$\mathbb{E}\left[\left(\delta_{\mathcal{B}}(Z)-\Theta\right)^{\top}(\delta(Z)-\delta_{\mathcal{B}}(Z))\right]$ $=\mathbb{E}\left[\left(\delta_{\mathcal{B}}(Z)-\mathbb{E}[\Theta\,|\,Z]\right)^{\top}(\delta(Z)-\delta_{\mathcal{B}}(Z))\right]$
$=\mathbb{E}\left[0^{\top}(\delta(Z)-\delta_{\mathcal{B}}(Z))\right]=0.$
In particular, we have shown
$\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]=R_{\mathcal{B}}+\textnormal{Var}\left(\delta(Z)-\delta_{\mathcal{B}}(Z)\right),$
which we regard as a decomposition of the objective of ($\mathcal{VCB}$) into the Bayes risk and a certain excess risk that we must minimize.
The preceding paragraph shows that the solution set to problem ($\mathcal{VCB}$) is equal to the solution set of the optimization problem
$\begin{cases}\textnormal{minimize}&\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta]\\ \textnormal{and}&\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta).\end{cases}$ (F.1)
Now we apply a “Gaussianization trick” whereby we show that problem (F.1) has the same optimal value as the following problem:
$\begin{cases}\textnormal{minimize}&\textnormal{Var}(D-E)\\ \textnormal{over}&\textnormal{jointly Gaussian }(D,E)\\ \textnormal{with}&\textnormal{marginal distribution }D\sim\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\Theta))\\ \textnormal{and}&\textnormal{marginal distribution }E\sim\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))\\ \end{cases}$ (F.2)
To see that (F.2) $\leq$ (F.1), it suffices to find, for each $\delta$ that is feasible for (F.1), some $(D,E)$ that is feasible for (F.2) such that $\textnormal{Var}(D-E)=\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))$. Of course, we can let
$\begin{pmatrix}D\\ E\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mathbb{E}[\Theta]\\ \mathbb{E}[\Theta]\end{pmatrix},\begin{pmatrix}\textnormal{Cov}(\Theta)&\textnormal{Cov}(\delta(Z),\delta_{\mathcal{B}}(Z))\\ \textnormal{Cov}(\delta_{\mathcal{B}}(Z),\delta(Z))&\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\end{pmatrix}\right)$ (F.3)
where $\textnormal{Cov}(X,Y):=\mathbb{E}[XY^{\top}]-\mathbb{E}[X]\mathbb{E}[Y]^{\top}$ denotes the cross-covariance between two random vectors $X$ and $Y$ of the same dimension. In this case $(D,E)$ and $(\delta(Z),\delta_{\mathcal{B}}(Z))$ have the same second-order structure, so we have $\textnormal{Var}(D-E)=\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))$.
To see that (F.1) $\leq$ (F.2), it suffices to show that the optimal value of (F.2) is achieved by a coupling which corresponds to some feasible $\delta$ for (F.1) (via (F.3)). When combined with the above, this will show that such a $\delta$ is optimal for (F.1).
Note that (F.2) is indeed the (squared) 2-Wasserstein distance between the two multivariate normal distributions $\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))$ and $\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\Theta))$. Indeed, recall [Dowson and Landau, 1982, Olkin and Pukelsheim, 1982] that the optimal value of (F.2) is exactly given by the Bures-Wasserstein distance
$\Pi^{2}(\textnormal{Cov}(\Theta),\textnormal{Cov}(\delta_{\mathcal{B}}(Z))),$
where $\Pi(\Sigma,\Sigma^{\prime}):=W_{2}(\mathcal{N}(0,\Sigma),\mathcal{N}(0,\Sigma^{\prime}))$, so, in particular we can use (PD) to define
$\delta_{\mathcal{VCB}}(z):=\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}(\delta_{\mathcal{B}}(z)-\mathbb{E}[\Theta])+\mathbb{E}[\Theta].$
[…]
$\textnormal{Var}(\delta_{\mathcal{VCB}}(Z)-\delta_{\mathcal{B}}(Z))$ $=\textnormal{Var}\left(\left(\mathbf{t}_{\textnormal{Cov}(\mathbb{E}[\Theta\,|\,Z])}^{\textnormal{Cov}(\Theta)}-I\right)\delta_{\mathcal{B}}(Z)\right)$
$=\textnormal{tr}\left(\left(\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}-I\right)\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\left(\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}-I\right)\right)$
$=\left\|\left(\mathbf{t}_{\textnormal{Cov}(\delta_{\mathcal{B}}(Z))}^{\textnormal{Cov}(\Theta)}-I\right)\Big{(}\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\Big{)}^{\nicefrac{{1}}{{2}}}\right\|_{2}^{2}$
$=\Pi^{2}\Big{(}\textnormal{Cov}(\Theta),\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\Big{)}.$
This completes the proof. For the last claim, note that assumptions (BC) and (ZC) allow us to apply García-Trillos and Sen [2024, Theorem 2.4]. ∎
Proof of Corollary 3.5.
We of course have $R_{\mathcal{B}}\leq R_{\mathcal{VCB}}$ by nature of imposing constraints, so it suffices to prove $R_{\mathcal{VCB}}\leq 2R_{\mathcal{B}}$ which is equivalent to $R_{\mathcal{VCB}}-R_{\mathcal{B}}\leq R_{\mathcal{B}}$. Since $\mathbb{E}[\delta_{\mathcal{B}}(Z)]=\mathbb{E}[\Theta]=:\mu$, we can use the tower property to compute:
$R_{\mathcal{B}}$ $=\mathbb{E}\left[\left\|\delta_{\mathcal{B}}(Z)-\Theta\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|\Theta-\mu\right\|^{2}\right]-\mathbb{E}\left[\left\|\delta_{\mathcal{B}}(Z)-\mu\right\|^{2}\right]$
$=\textnormal{tr}(\textnormal{Cov}(\Theta))-\textnormal{tr}(\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))$
$=\textnormal{tr}(\textnormal{Cov}(\Theta)-\textnormal{Cov}(\delta_{\mathcal{B}}(Z))).$
Now recall $\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\preceq\textnormal{Cov}(\Theta)$ from (1.2), so we have shown
$R_{\mathcal{B}}=\left\|\textnormal{Cov}(\Theta)-\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\right\|_{1},$
where $\|\cdot\|_{1}$ denotes the trace norm (also called the nuclear norm or the Schatten 1-norm), defined as $\|A\|_{1}:=\textnormal{tr}(|A|)$ for a real symmetric matrix $A$, where $|A|$ is its matrix absolute value. Now note that Theorem 3.3 implies
$R_{\mathcal{VCB}}-R_{\mathcal{B}}=\Pi^{2}\left(\textnormal{Cov}(\Theta),\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\right).$
Thus, it suffices to show that we have
$\Pi^{2}\left(\textnormal{Cov}(\Theta),\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\right)\leq\left\|\textnormal{Cov}(\Theta)-\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\right\|_{1}.$
To show this, we let $\|\cdot\|_{2}$ denotes the Frobenius norm (also called the Schatten 2-norm), defined as $\|A\|_{2}:=\sqrt{\textnormal{tr}(A^{2})}$ for a real symmetric matrix $A$. Then we note that we in fact have $\Pi^{2}(A,B)\leq\|A^{\nicefrac{{1}}{{2}}}-B^{\nicefrac{{1}}{{2}}}\|_{2}^{2}\leq\|A-B\|_{1}$ for all $A,B\in\mathbb{K}(m)$ as a consequence of the Procrustes representation of the Bures-Wasserstein metric [Santoro and Panaretos, 2023, Lemma 1] and the Powers-Størmer inequality [Powers and Størmer, 1970, Lemma 4.1], respectively. (See also [Santoro and Panaretos, 2023, Lemma 1].) ∎
Proof of Theorem C.3.
First let us show that the integrand of (C.2) is well-defined for every $\pi\in\Gamma(F;G,\Psi)$. To see this, write $H$ for the second marginal of $\pi$, note that
$c_{G}(z,\eta)=\|\eta-\delta_{\mathcal{B}}(z)\|^{2}\leq 2\|\eta\|^{2}+2\|\delta_{\mathcal{B}}(z)\|^{2}.$
Then, use the contraction property of conditional expectation to bound
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\textnormal{d}\pi(z,\eta)$ $\leq 2\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}H(\eta)+2\int_{\mathbb{R}^{d}}\|\delta_{\mathcal{B}}(z)\|^{2}\textnormal{d}F(z)$
$\leq 2\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}H(\eta)+2\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}G(\eta).$
By (QG), we may get a function $\psi\in\Psi$ and some $C,R>0$ such that we have $\|\eta\|^{2}\leq C\psi(\eta)$ for all $\eta\in\mathbb{R}^{m}$ with $\|\eta\|>R$. Then we can bound the right side via
$\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}H(\eta)$ $=\int_{B_{r}(0)}\|\eta\|^{2}\textnormal{d}H(\eta)+\int_{\mathbb{R}^{m}\setminus B_{r}(0)}\|\eta\|^{2}\textnormal{d}H(\eta)$
$\leq\int_{B_{r}(0)}\|\eta\|^{2}\textnormal{d}H(\eta)+C\int_{\mathbb{R}^{m}\setminus B_{r}(0)}\psi(\eta)\textnormal{d}H(\eta)$
$\leq R+C\int_{\mathbb{R}^{m}\setminus B_{r}(0)}\psi(\eta)\textnormal{d}G(\eta)$
$\leq R+C\int_{\mathbb{R}^{m}}\psi\textnormal{d}G,$
and similar for $G$. Since the right side above is finite by the assumption $\psi\in\Psi\subseteq L^{1}(G)$, it follows that the integral is well-defined for each feasible $\pi$.
Next let us show that (C.2) admits solutions. By (C), there exists a Borel set $A\subseteq\mathbb{R}^{d}$ with $F(A)=1$ such that $c_{G}:A\times\mathbb{R}^{m}\to[0,\infty)$ is continuous. Since
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\textnormal{d}\pi(z,\eta)=\int_{A\times\mathbb{R}^{m}}c_{G}(z,\eta)\textnormal{d}\pi(z,\eta),$
for all $\pi\in\Gamma(F;G)$, it follows from the Portmanteau lemma that the objective of (C.2) is weakly lower semi-continuous. Since a lower semi-continuous function on a compact set must achieve a minimizer, it only remains to show that the feasible set $\Gamma(F;G,\Psi)$ is weakly compact. To do this, note that we can write $\Gamma(F;G,\Psi)=\bigcup_{H\in\mathcal{H}}\Gamma(F;H)$ for
:=$\mathcal{H}:=\left\{H\in\mathcal{P}(\mathbb{R}^{m}):\int_{\mathbb{R}^{m}}\phi_{\ell}(\eta)\textnormal{d}H(\eta)=\int_{\mathbb{R}^{m}}\phi_{\ell}(\eta)\textnormal{d}G(\eta)\textnormal{ for all }1\leq\ell\leq k\right\},$
and it follows from [Villani, 2009, Lemma 4.3] that $\Gamma(F;G,\Psi)$ is weakly compact in $\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ as soon as $\mathcal{H}$ is weakly compact in $\mathcal{P}(\mathbb{R}^{m})$. To show the latter, it suffices by Prokhorov’s theorem to show that $\mathcal{H}$ is tight. To do this, use (QG) to get (as above) some $\psi\in\Psi$ and $C,R>0$ with $\|\eta\|^{2}\leq C\psi(\eta)$ for all $\eta\in\mathbb{R}^{m}$ with $\|\eta\|>R$. In particular, the compact sets $K_{r}:=\{\eta\in\mathbb{R}^{m}:\psi(\eta)\leq r\}$ for $r\geq 0$ satisfy, by Markov’s inequality:
$H(\mathbb{R}^{m}\setminus K_{r})\leq\frac{1}{r}\int_{\mathbb{R}^{m}}\psi(\eta)\textnormal{d}H(\eta)=\frac{1}{r}\int_{\mathbb{R}^{m}}\psi(\eta)\textnormal{d}G(\eta)$
for all $H\in\mathcal{H}$ and $r\geq 0$, hence
$\sup_{H\in\mathcal{H}}H(\mathbb{R}^{m}\setminus K_{r})\leq\frac{1}{r}\int_{\mathbb{R}^{m}}\psi(\eta)\textnormal{d}G(\eta)\to 0$
as $r\to\infty$. This proves that $\mathcal{H}$ is tight, and hence guarantees the existence of a solution to (C.2).
Now we finish the proof, by applying the result in García-Trillos and Sen [2024, Theorem 2.4 and Remark 4.2] which shows that, under assumptions (BC) and (QG), any solution to (C.2) must be induced by a map. Indeed, it suffices to show that (C.2) admits a unique solution, so, towards a contradiction, assume that there were distinct solutions $\pi_{1},\pi_{2}$ to (C.2). Since the objective of (C.2) is affine, it follows that $\frac{1}{2}(\pi_{1}+\pi_{2})$ is also a solution. However, if $\pi_{1},\pi_{2}$ are induced by maps, then $\frac{1}{2}(\pi_{1}+\pi_{2})$ cannot be induced by a map. This is a contradiction, hence the proof is complete. ∎
Appendix G Proofs from Section 4
This section contains the proofs of the main results in Section 4 on EB denoising.
G.1 Proofs from Subsection 4.1
We begin with the proofs of the results from Subsection 4.1 on variance-constrained EB denoising. In particular, we will prove Theorem 4.3 which provides a rate of convergence for the denoiser $\hat{\delta}_{\mathcal{VCB}}$ given in Algorithm 1. Adopting the notation from Subsection 4.1, we begin with the following estimate.
Lemma G.1.
Under (FM) and (EBQ), we have
$\|\hat{M}-M\|_{\textnormal{F}}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)$
as $n\to\infty$.
Proof.
We begin by proving the following deterministic result: There exists a universal constant $C>0$ such that, for any $x_{1},\ldots,x_{n}$ and $y_{1},\ldots,y_{n}$ in $\mathbb{R}^{m}$, we have
$\left\|\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\top}\right\|_{\textnormal{F}}\leq CD(D+\tilde{D}),$ (G.1)
where we have defined $\bar{x}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and $\bar{y}:=\frac{1}{n}\sum_{i=1}^{n}y_{i}$, as well as
$D:=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-y_{i}\|^{2}}$
and
$\tilde{D}:=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|y_{i}-\bar{y}\|^{2}}.$
To see this, note that for all $u,v\in\mathbb{R}^{m}$ we have the bound $\|uu^{\top}-vv^{\top}\|_{\textnormal{F}}\leq\|u-v\|(\|u\|+\|v\|)$, hence we get
$\left\|\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\top}\right\|_{\textnormal{F}}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left\|(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}-(y_{i}-\bar{y})(y_{i}-\bar{y})^{\top}\right\|_{\textnormal{F}}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-y_{i}\|(\|x_{i}-\bar{x}\|+\|y_{i}-\bar{y}\|).$
By applying Cauchy-Schwarz to the right side and then squaring both sides of the resulting inequality, the display above yields
$\left\|\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\top}\right\|_{\textnormal{F}}^{2}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-y_{i}\|^{2}\cdot\frac{2}{n}\sum_{i=1}^{n}\left(\|x_{i}-\bar{x}\|^{2}+\|y_{i}-\bar{y}\|^{2}\right).$
In particular, it suffices to show that there is some universal constant $C>0$ such that we have
$\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-\bar{x}\|^{2}\leq C\left(\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-y_{i}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}\|y_{i}-\bar{y}\|^{2}\right)$
To do this, we use the elementary inequality
$\|x_{i}-\bar{x}\|^{2}\leq 3\left(\|x_{i}-y_{i}\|^{2}+\|y_{i}-\bar{y}\|^{2}+\|\bar{y}-\bar{x}\|^{2}\right)$
and the observation
$\|\bar{x}-\bar{y}\|^{2}\leq\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-y_{i}\|^{2}$
to conclude.
Now we return to the main result. Set
$x_{i}:=\hat{\delta}_{\mathcal{B}}(Z_{i})\qquad\mbox{and}\qquad y_{i}:=\delta_{\mathcal{B}}(Z_{i}),$
in (G.1) an note that $\hat{M}$ is the sample covariance matrix of $\{x_{i}\}_{i=1}^{n}$ and $M$ is the population covariance matrix of $\{y_{i}\}_{i=1}^{n}$. Letting $\bar{M}$ denote the sample covariance matrix
$\bar{M}:=\frac{1}{n}\sum_{i=1}^{n}y_{i}y_{i}^{\top}-\bar{y}\bar{y}^{\top},$
we note that by the triangle inequality we have
$\|\hat{M}-M\|_{\textnormal{F}}^{2}\leq 2\|\hat{M}-\bar{M}\|_{\textnormal{F}}^{2}+2\|\bar{M}-M\|_{\textnormal{F}}^{2}.$ (G.2)
Since $y_{1},\ldots,y_{n}$ are i.i.d., we know from classical asymptotic statistics [Durrett, 2019, Theorem 2.5.12] that the second term is $O_{\mathbb{P}}(n^{-2(q-2)/q})$ provided that $\mathbb{E}[\|\delta_{\mathcal{B}}(Z)\|^{q}]<\infty$; to see this, simply use the contraction property of conditional expectation to compute:
$\mathbb{E}\left[\|\delta_{\mathcal{B}}(Z)\|^{q}\right]=\mathbb{E}\left[\|\mathbb{E}[\Theta\,|\,Z]\|^{q}\right]\leq\mathbb{E}\left[\|\Theta\|^{q}\right]<\infty,$
where the finiteness holds by (FM). For the first term, note by (G.1) that it suffices to show $D^{2}=O_{\mathbb{P}}(\alpha_{n})$ and $\tilde{D}^{2}=O_{\mathbb{P}}(1)$; the latter follows from classical asymptotic statistics, as above, and the former holds by assumption (EBQ). This finishes the proof. ∎
Now we can prove the main result.
Proof of Theorem 4.3.
Since all elements of the problem are equivariant under translation, we may assume $\mu:=\mathbb{E}[Z]=\mathbb{E}[\Theta]$ is $\mu=0$. Then write $\hat{\mathbf{t}}=\mathbf{t}_{\hat{M}}^{\hat{A}}$ so that we can expand
$\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})$ $=\left(\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right)\hat{\delta}_{\mathcal{B}}(Z_{i})+\mathbf{t}_{M}^{A}\left(\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right)+(I-\mathbf{t}_{\hat{M}}^{\hat{A}})\hat{\mu},$
and let us use this to bound
$\begin{split}\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})\right\|_{2}^{2}&\leq\left\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}\frac{3}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\\ &+\left\|\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}\frac{3}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\\ &+3\left\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\right\|_{2\to 2}^{2}\|\hat{\mu}\|_{2}^{2},\end{split}$ (G.3)
where $\|\cdot\|_{2\to 2}$ denotes the operator norm of a linear map from $(\mathbb{R}^{m},\|\cdot\|)$ to itself. In order to complete the proof, it suffices to show that the three terms on the right side are all $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ as $n\to\infty$. We do this in reverse order.
For the third term of (G.3), it suffices to show first $\mathbb{P}(\hat{M}\preceq\hat{A})\to 1$ and then $n\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\|_{2\to 2}^{2}\|\hat{\mu}\|_{2}^{2}=O_{\mathbb{P}}(1)$ on $\{\hat{M}\preceq\hat{A}\}$; this is because $2(q-2)/q\leq 1$ for all $2<q\leq 4$, hence $O_{\mathbb{P}}(n^{-1})\leq O_{\mathbb{P}}(n^{-2(q-2)/q})$. The former follows since $\{(M,A)\in\mathbb{K}(m):M\not\prec A\}$ is an open set containing $(M,A)$ (recall (1.2)), and we have $\hat{M}\to M$ by Lemma G.1, as well as $\hat{A}\to A$ by combining (EC) with $\hat{S}\to S$ that follows from (FM) and the strong law of large numbers. To see the latter, note that $\hat{M}\preceq\hat{A}$ implies $\mathbf{t}_{\hat{M}}^{\hat{A}}\preceq I$ hence $\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\|_{2\to 2}\leq 1$. Moreover, we have $\mathbb{E}[\|Z\|^{2}]=\textnormal{tr}(A)+\textnormal{tr}(\Sigma)<\infty$ by (M) and (FM), so classical asymptotic statistics yields $\|\hat{\mu}\|_{2}^{2}=O_{\mathbb{P}}(n^{-1})$.
For the second term, we simply note that the first factor is non-random and that the second factor is $O_{\mathbb{P}}(\alpha_{n})$ by assumption (EBQ).
Lastly, we consider the first term of (G.3). Indeed, we note that it suffices to show
$\left\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)$ (G.4)
and
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}=O_{\mathbb{P}}(1)$ (G.5)
as $n\to\infty$, and we do this in two separate steps.
To see (G.4), we use the triangle inequality to get
$\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\|_{2\to 2}\leq\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{\hat{A}}\|_{2\to 2}+\|\mathbf{t}_{M}^{\hat{A}}-\mathbf{t}_{M}^{A}\|_{2\to 2}.$
Of course, all norms are equivalent up to constants which may depend only on the dimension, so it suffices to show that both
$\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{\hat{A}}\|_{\textnormal{F}}\qquad\textnormal{and}\qquad\|\mathbf{t}_{M}^{\hat{A}}-\mathbf{t}_{M}^{A}\|_{\textnormal{F}}$
are $O_{\mathbb{P}}(\alpha_{n}\vee n^{-1})$ as $n\to\infty$. To do this, we define $\phi^{B}:\mathbb{K}(m)\to\mathbb{S}(m)$ via $\phi^{B}(A):=\mathbf{t}_{A}^{B}$, and note from Kroshnin et al. [2021, Lemma A.2] that $\phi^{B}$ is continuously differentiable and its Jacobian $\nabla_{A}\phi^{B}$ is jointly continuous in $(A,B)$. At the same time, the mean value theorem yields
$\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{\hat{A}}\|_{\textnormal{F}}\leq\sup_{0\leq t\leq 1}\|\nabla_{(1-t)\hat{M}+tM}\phi^{\hat{A}}\|_{\textnormal{F}\to\textnormal{F}}\cdot\|\hat{M}-M\|_{\textnormal{F}},$
where $\|\cdot\|_{\textnormal{F}\to\textnormal{F}}$ denotes the operator norm of a linear map from $(\mathbb{R}^{m\times m},\|\cdot\|_{\textnormal{F}})$ to itself. Now fix $r>0$, and note that the ball $B_{r}^{\textnormal{F}}(B):=\{A\in\mathbb{K}(m):\|A-B\|_{\textnormal{F}}\leq r\}$ is compact and convex for all $B\in\mathbb{R}^{m\times m}$. In particular, we have
$C_{1}:=\sup_{\tilde{M}\in B^{\textnormal{F}}_{r}(M),\tilde{A}\in B^{\textnormal{F}}_{r}(A)}\|\nabla_{\tilde{M}}\phi^{\tilde{A}}\|_{\textnormal{F}\to\textnormal{F}}<\infty,$
and
$\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{\hat{A}}\|_{\textnormal{F}}\leq C_{1}\|\hat{M}-M\|_{\textnormal{F}},$
on the event $\{\|\hat{M}-M\|_{\textnormal{F}}\leq r,\|\hat{A}-A\|_{\textnormal{F}}\leq r\}$. Additionally, note that we may define the map $\phi_{A}:\mathbb{K}(m)\to\mathbb{S}(m)$ via $\phi_{A}(B):=\mathbf{t}_{A}^{B}$, and of course this is related to $\phi^{B}$ by the identity $\phi_{A}=(\phi^{A})^{-1}$. Since inversion is differentiable on $\mathbb{K}(m)$, this implies that $\phi_{A}$ is also continuously differentiable. In particular, we have
$C_{2}:=\sup_{\tilde{A}\in B^{\textnormal{F}}_{r}(A)}\|\nabla_{\tilde{A}}\phi_{M}\|_{\textnormal{F}\to\textnormal{F}}<\infty,$
and
$\|\mathbf{t}_{M}^{\hat{A}}-\mathbf{t}_{M}^{A}\|_{\textnormal{F}}\leq C_{2}\|\hat{A}-A\|_{\textnormal{F}},$
on the event $\{\|\hat{A}-A\|_{\textnormal{F}}\leq r\}$. Since we have $\mathbb{P}(\|\hat{M}-M\|_{\textnormal{F}}\leq r,\|\hat{A}-A\|_{\textnormal{F}}\leq r)\to 1$ by (EC) and Lemma G.1, it suffices to show that all three of
$\|\hat{S}-S\|_{\textnormal{F}},\qquad\|\hat{\Sigma}-\Sigma\|_{\textnormal{F}},\qquad\textnormal{ and }\qquad\|\hat{M}-M\|_{\textnormal{F}}$
are $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ as $n\to\infty$. As we have already seen, the first is $O_{\mathbb{P}}(n^{-2(q-2)/q})$ from classical asymptotic statistics, the second is $O_{\mathbb{P}}(n^{-1})\leq O_{\mathbb{P}}(n^{-2(q-2)/q})$ by (EC), and the third is $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ by Lemma G.1. This establishes (G.4).
To see (G.5), we simply bound:
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\leq\frac{2}{n}\sum_{i=1}^{n}\left\|\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}+\frac{2}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}.$ (G.6)
Lastly, note that the first term is $O_{\mathbb{P}}(1)$ by classical asymptotic statistics and (FM), and the second term is $O_{\mathbb{P}}(\alpha_{n})$ by another application of (EBQ). This shows that the sum is $O_{\mathbb{P}}(1)$, which establishes (G.5). This finishes the proof. ∎
G.2 Proofs from Subsection 4.2
Next we turn to the results of Subsection 4.2 on EB denoising with distribution constraints. In particular, we will prove Theorem 4.4 which provides a rate of convergence of the denoiser $\hat{\delta}_{\mathcal{DCB}}$ given in Algorithm 2.
Proof of Theorem 4.4.
We begin with a useful estimate, which is based on the proof of Deb et al. [2021, Theorem 2.1] and Slawski and Sen [2024, Theorem 7]. To derive it, recall from Theorem 3.8 that $\delta_{\mathcal{DCB}}$ must be of the form $\nabla\phi\circ\delta_{\mathcal{B}}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}$. Consequently, we have $(\nabla\phi^{\ast})(\delta_{\mathcal{DCB}}(z))=\delta_{\mathcal{B}}(z)$ holding for $F$-almost all $z\in\mathbb{R}^{d}$, where $\phi^{\ast}$ is the Fenchel-Legendre dual of $\phi$. Also, let us define
$\bar{G}_{n}:=(\delta_{\mathcal{DCB}})_{\#}\bar{F}_{n}\qquad\textnormal{and}\qquad\bar{\pi}_{n}:=(\textnormal{id},\delta_{\mathcal{DCB}})_{\#}\bar{F}_{n}$
for each $n\in\mathbb{N}$ (recall that $\bar{F}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{Z_{i}}$). Then, we can use the smoothness from (R) to compute:
$\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\,\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\,\textnormal{d}\bar{G}_{n}(\eta)$
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\,\textnormal{d}\hat{\pi}_{n}(\eta\,|\,z)\,\textnormal{d}\bar{F}_{n}(z)-\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\,\textnormal{d}\bar{G}_{n}(\eta)$
$\geq\int_{\mathbb{R}^{d}}\phi^{\ast}(\hat{\delta}_{\mathcal{DCB}}(z))\,\textnormal{d}\bar{F}_{n}(z)-\int_{\mathbb{R}^{d}}\phi^{\ast}(\delta_{\mathcal{DCB}}(z))\,\textnormal{d}\bar{F}_{n}(z)$
$=\int_{\mathbb{R}^{d}}\left(\phi^{\ast}(\hat{\delta}_{\mathcal{DCB}}(z))-\phi^{\ast}(\delta_{\mathcal{DCB}}(z))\right)\textnormal{d}\bar{F}_{n}(z)$
$\geq\int_{\mathbb{R}^{d}}\left\langle(\nabla\phi^{\ast})(\delta_{\mathcal{DCB}}(z)),\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)$
$\qquad\qquad\qquad+\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)$
$=\int_{\mathbb{R}^{d}}\left\langle(\delta_{\mathcal{B}}(z),\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)$
$\qquad\qquad\qquad+\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)$
Here, the first inequality follows from Jensen and the second inequality follows from assumption (R). Rearranging the above, we have shown
$\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\,\textnormal{d}\bar{F}_{n}(z)$$\leq$ $\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\textnormal{d}\bar{G}_{n}(\eta)$ (G.7)
$\;\;-\int_{\mathbb{R}^{d}}\left\langle(\delta_{\mathcal{B}}(z),\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z).\qquad$
Now we make two computations. First, we get
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}$ $=$ $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\|\hat{\delta}_{\mathcal{B}}(z)-\eta\|^{2}\textnormal{d}\hat{\pi}_{n}(z,\eta)$ (G.8)
$=$ $\int_{\mathbb{R}^{d}}\|\hat{\delta}_{\mathcal{B}}(z)\|^{2}\textnormal{d}\bar{F}_{n}(z)+\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}\hat{G}_{n}(\eta)-2\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{m}}\left\langle\eta,\hat{\delta}_{\mathcal{B}}(z)\right\rangle\textnormal{d}\hat{\pi}_{n}(\eta\,|\,z)\textnormal{d}\bar{F}_{n}(z)$
$=$ $\int_{\mathbb{R}^{d}}\|\hat{\delta}_{\mathcal{B}}(z)\|^{2}\textnormal{d}\bar{F}_{n}(z)+\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}\hat{G}_{n}(\eta)-2\int_{\mathbb{R}^{d}}\left\langle\hat{\delta}_{\mathcal{DCB}}(z),\hat{\delta}_{\mathcal{B}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)$
from the fact that we have $\hat{\delta}_{\mathcal{DCB}}(z)=\int\eta\,\textnormal{d}\hat{\pi}_{n}(\eta\,|\,z)$ by construction. Second, we get
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}$ $=$ $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\|\delta_{\mathcal{B}}(z)-\eta\|^{2}\textnormal{d}\bar{\pi}_{n}(z,\eta)$ (G.9)
$=$ $\int_{\mathbb{R}^{d}}\|\delta_{\mathcal{B}}(z)\|^{2}\textnormal{d}\bar{F}_{n}(z)+\int_{\mathbb{R}^{m}}\|\eta\|^{2}\textnormal{d}\bar{G}_{n}(\eta)-2\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z),\quad\qquad$
from the fact that we have ${\delta}_{\mathcal{DCB}}(z)=\int\eta\,\textnormal{d}\bar{\pi}_{n}(\eta\,|\,z)$ holding $\bar{F}_{n}$ almost surely, which requires a small argument: defining $\bar{\delta}_{\mathcal{DCB}}(z):=\int\eta\,\textnormal{d}\bar{\pi}_{n}(\eta\,|\,z)$, we may use the contraction property of conditional expectation to compute:
$\int_{\mathbb{R}^{d}}\|\bar{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{F}_{n}(z)$
$=\int_{\mathbb{R}^{d}}\|\bar{\delta}_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{F}_{n}(z)+\int_{\mathbb{R}^{d}}\|\delta_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{F}_{n}(z)-2\int_{\mathbb{R}^{d}}\langle\delta_{\mathcal{DCB}}(z),\bar{\delta}_{\mathcal{DCB}}(z)\rangle\textnormal{d}\bar{F}_{n}(z)$
$=\int_{\mathbb{R}^{d}}\|\bar{\delta}_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{\pi}_{n}(z,\eta)+\int_{\mathbb{R}^{d}}\|\delta_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{\pi}_{n}(z,\eta)-2\int_{\mathbb{R}^{d}}\langle\delta_{\mathcal{DCB}}(z),\bar{\delta}_{\mathcal{DCB}}(z)\rangle\textnormal{d}\bar{\pi}_{n}(z,\eta)$
$\leq\int_{\mathbb{R}^{d}}\|\eta\|^{2}\textnormal{d}\bar{\pi}_{n}(z)+\int_{\mathbb{R}^{d}}\|\delta_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{\pi}_{n}(z)-2\int_{\mathbb{R}^{d}}\langle\delta_{\mathcal{DCB}}(z),\eta\rangle\textnormal{d}\bar{\pi}_{n}(z)$
$=\int_{\mathbb{R}^{d}}\|\eta-\delta_{\mathcal{DCB}}(z)\|^{2}\textnormal{d}\bar{\pi}_{n}(z)$
$=0.$
Thus, equations (G.9) and (G.8) are both true. By subtracting (G.9) from (G.8), and plugging the result into (G.7) and rearranging, we conclude:
$\begin{split}\frac{1}{2L}&\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\leq\int_{\mathbb{R}^{m}}\left(\phi^{\ast}(\eta)-\frac{1}{2}\|\eta\|^{2}\right)\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\left(\phi^{\ast}(\eta)-\frac{1}{2}\|\eta\|^{2}\right)\textnormal{d}\bar{G}_{n}(\eta)\\ &\qquad-2\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)+\int\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\int\big{\|}\hat{\delta}_{\mathcal{B}}(z)\big{\|}^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\qquad+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}\right)+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}\right).\end{split}$ (G.10)
Next, let $\rho_{n}$ be an optimal coupling of $\bar{G}_{n}$ and $\hat{G}_{n}$, and define $\lambda_{n}:=(\nabla\phi^{\ast},\textnormal{id})_{\#}\rho_{n}$, so that $\lambda_{n}$ is a coupling (not necessarily optimal) of $(\delta_{\mathcal{B}})_{\#}\bar{F}_{n}$ and $\hat{G}_{n}$. Then we can compute:
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\eta-\hat{\eta}\|^{2}\textnormal{d}\lambda_{n}(\eta,\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\hat{\eta}\|^{2}\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})+\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\bar{\eta}-\hat{\eta}\|^{2}\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &\qquad+2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta}).\end{split}$ (G.11)
Note that the first term on the right side is exactly
$\int_{\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})=\int_{\mathbb{R}^{m}}\|\delta_{\mathcal{B}}(z)-\nabla\phi(\delta_{\mathcal{B}}(z))\|^{2}\textnormal{d}\bar{F}_{n}(z)=W_{2}^{2}((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n})$
because $\nabla\phi$ is an optimal transport map, and that the second term is $W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})$ by definition. Therefore, (G.11) becomes
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\eta-\hat{\eta}\|^{2}\textnormal{d}\lambda_{n}(\eta,\hat{\eta})\\ &=W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right)+W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})\\ &\qquad+2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta}).\end{split}$ (G.12)
For the cross term, expand the square and use the strong convexity from (R) to get
$\begin{split}&2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta}),\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &\leq 2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left(\phi^{\ast}(\bar{\eta})-\phi^{\ast}(\hat{\eta})+\frac{1}{2\lambda}\|\bar{\eta}-\hat{\eta}\|^{2}\right)\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=2\int_{\mathbb{R}^{m}}\phi^{\ast}(\bar{\eta})\textnormal{d}\bar{G}_{n}(\bar{\eta})-2\int_{\mathbb{R}^{m}}\phi^{\ast}(\hat{\eta})\textnormal{d}\hat{G}_{n}(\bar{\eta})+\frac{1}{\lambda}W_{2}^{2}(\hat{G}_{n},\hat{G}_{n})\end{split}$ (G.13)
and
$\begin{split}&2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left(\|\hat{\eta}\|^{2}-\|\hat{\eta}-\bar{\eta}\|^{2}-\|\bar{\eta}\|^{2}\right)\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}}\|\hat{\eta}\|^{2}\textnormal{d}\hat{G}_{n}(\hat{\eta})-\int_{\mathbb{R}^{m}}\|\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})-W_{2}^{2}(\bar{G}_{n},\hat{G}_{n}),\end{split}$ (G.14)
and then plug in (G.13) and (G.14) into (G.12), yielding
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right)+\frac{1}{\lambda}W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})\\ &\quad+\int_{\mathbb{R}^{m}}\left(2\phi^{\ast}(\bar{\eta})-\|\bar{\eta}\|^{2}\right)\textnormal{d}\bar{G}_{n}(\bar{\eta})-\int_{\mathbb{R}^{m}}\left(2\phi^{\ast}(\hat{\eta})-\|\hat{\eta}\|^{2}\right)\textnormal{d}\hat{G}_{n}(\bar{\eta}).\end{split}$ (G.15)
Since
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}=W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)\qquad\textnormal{ and }\qquad\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}=W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right),$
we may plug in (G.15) into (G.10) to get:
$\begin{split}\frac{1}{2L}&\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\leq\;\frac{1}{2\lambda}W_{2}^{2}(\hat{G}_{n},\bar{G}_{n})-2\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)\\ &+\int_{\mathbb{R}^{d}}\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}\right).\\ \end{split}$
Now, expand the squares in the definitions of the last two terms to get:
$\begin{split}\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)&\leq\frac{1}{2\lambda}W_{2}^{2}(\hat{G}_{n},\bar{G}_{n})\\ &-\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)\\ &+\frac{1}{2}\int_{\mathbb{R}^{d}}\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\frac{1}{2}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z).\end{split}$ (G.16)
It is useful to compare (G.16) with Slawski and Sen [2024, Theorem 7] which derives a similar bound, but with only the first term; the additional terms are due to the fact that $\hat{\delta}_{\mathcal{DCB}}$ comes from an transport problem whose cost function has been estimated.
In the rest of the proof, we will show that each of the terms on the right side of (G.16) is $O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\beta_{n}\vee\gamma_{n})$. The first term is the easiest to handle since by the triangle inequality we can bound
$W_{2}^{2}(\hat{G}_{n},\bar{G}_{n})\leq 2W_{2}^{2}(\hat{G}_{n},G)+2W_{2}^{2}(\bar{G}_{n},G),$
and note by assumption (D) that these terms on the right side are $O_{\mathbb{P}}(\beta_{n})$ and $O_{\mathbb{P}}(\gamma_{n})$, respectively. For the second term, use Cauchy-Schwarz to bound
$\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)$
$\leq\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{DCB}}(Z_{i})\|^{2}}.$
and note that the first factor is $O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}})$ by (EBQ), and the second factor is $O_{\mathbb{P}}(1)$ because of (2M) and the law of large numbers. For the third term, recall the elementary inequality
$\left|\|x\|^{2}-\|y\|^{2}\right|\leq\|x-y\|^{2}+2\|x\|\cdot\|x-y\|$
for all $x,y\in\mathbb{R}^{m}$. Taking $x=\delta_{\mathcal{B}}(Z_{i})$ and $y=\hat{\delta}_{\mathcal{B}}(Z_{i})$, summing, and applying Cauchy-Schwarz yields
$\int_{\mathbb{R}^{d}}\|\delta_{\mathcal{B}}(Z_{i})\|^{2}\textnormal{d}\bar{F}_{n}(z)-\int_{\mathbb{R}^{d}}\|\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}\textnormal{d}\bar{F}_{n}(z)$
$=\sum_{i=1}^{n}\left(\|\delta_{\mathcal{B}}(Z_{i})\|^{2}-\|\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}\right)$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left(\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}+2\|\delta_{\mathcal{B}}(Z_{i})\|\cdot\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}+\frac{2}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})\|\cdot\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}+2\cdot\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})\|^{2}}.$
Note that the first term on the right is $O_{\mathbb{P}}(\alpha_{n})$ by assumption (EBQ). Within the second term, the first factor is $O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}})$ by assumption (EBQ), and the second factor is $O_{\mathbb{P}}(1)$ by assumption (2M) and the law of large numbers. ∎
G.3 Proofs from Subsection 4.3
Lastly, we give the proof of Theorem C.5 on the rate of convergence for the general-constrained EB denoiser $\hat{\delta}_{\mathcal{GCB}}$ from Algorithm 3.
An important step in our proof is a following technical result which states that limiting feasible couplings can be approximated by prelimiting feasible couplings. To state it, let us write $\Gamma_{K}(F;G,\Psi)$ for the set of $\pi\in\Gamma(F;G,\Psi)$ satisfying $\textnormal{supp}(\pi(\mathbb{R}^{d}\times\textnormal{d}\eta))\subseteq K$, and similar for $\Gamma_{K}(\bar{F}_{n};\hat{G}_{n},\Psi)$.
Lemma G.2.
Under assumptions (EC), (CS), (INT) and (CC), for every $\pi\in\Gamma_{K}(F;G,\Psi)$ there exists $\pi_{n}\in\Gamma_{K}(\bar{F}_{n};\hat{G}_{n},\Psi)$ for each $n\in\mathbb{N}$ such that $W_{2}(\pi_{n},\pi)=o_{\mathbb{P}}(1)$.
Proof.
First, take $\pi\in\Gamma_{K}(F;G,\Psi)$ and get $H_{0}\in\mathcal{P}(\mathbb{R}^{m})$ such that we have $\pi\in\Gamma(F;H_{0})$, and let construct a suitable exponential family from these pieces; this will require some basic notions which can be found in Wainwright and Jordan [2008, Chapter 3]. That is, let us define $A:\mathbb{R}^{k}\to\mathbb{R}\cup\{\infty\}$ via
$A(\alpha):=\int_{\mathbb{R}^{m}}\exp\left(\sum_{\ell=1}^{k}\alpha_{\ell}\psi_{\ell}(\eta)\right)\textnormal{d}H_{0}(\eta),$
and, for each $\alpha\in D(H_{0},\Phi)$ we define the probability measure $H_{\alpha}\in\mathcal{P}(\mathbb{R}^{m})$ via its Radon-Nikodym derivative
$\frac{\textnormal{d}H_{\alpha}}{\textnormal{d}H_{0}}(\eta)=\exp\left(\sum_{\ell=1}^{k}\alpha_{\ell}\psi_{\ell}(\eta)-A(\alpha)\right).$
It follows that $\{H_{\alpha}\}_{\alpha}$ is an exponential family; in fact, the linear independence of $\psi_{1},\ldots,\psi_{k}$ implies that $\{H_{\alpha}\}_{\alpha}$ is minimal (but not necessarily complete). Also, observe that $H_{0}$ indeed corresponds to the member of this family with $\alpha=0$. By assumption (CC) and (INT), we have
$\begin{pmatrix}\int_{\mathbb{R}^{m}}\psi_{1}\textnormal{d}\hat{G}_{n}\\ \vdots\\ \int_{\mathbb{R}^{m}}\psi_{k}\textnormal{d}\hat{G}_{n}\end{pmatrix}\to\begin{pmatrix}\int_{\mathbb{R}^{m}}\psi_{1}\textnormal{d}G\\ \vdots\\ \int_{\mathbb{R}^{m}}\psi_{k}\textnormal{d}G\end{pmatrix}\in\mathcal{M}^{\circ}(\Psi)$
as $n\to\infty$, so it follows that we have
$\begin{pmatrix}\int_{\mathbb{R}^{d}}\psi_{1}\textnormal{d}\hat{G}_{n}\\ \vdots\\ \int_{\mathbb{R}^{d}}\psi_{k}\textnormal{d}\hat{G}_{n}\end{pmatrix}\in\mathcal{M}^{\circ}(\Psi)$
for all sufficiently large $n\in\mathbb{N}$. This implies that, for such $n\in\mathbb{N}$, the parameter
$\alpha_{n}:=\nabla A^{\ast}\left(\begin{pmatrix}\int_{\mathbb{R}^{d}}\psi_{1}\textnormal{d}\hat{G}_{n}\\ \vdots\\ \int_{\mathbb{R}^{d}}\psi_{k}\textnormal{d}\hat{G}_{n}\end{pmatrix}\right)$
satisfies
$\begin{pmatrix}\int_{\mathbb{R}^{d}}\psi_{1}\textnormal{d}H_{\alpha_{n}}\\ \vdots\\ \int_{\mathbb{R}^{d}}\psi_{k}\textnormal{d}H_{\alpha_{n}}\end{pmatrix}=\begin{pmatrix}\int_{\mathbb{R}^{d}}\psi_{1}\textnormal{d}\hat{G}_{n}\\ \vdots\\ \int_{\mathbb{R}^{d}}\psi_{k}\textnormal{d}\hat{G}_{n}\end{pmatrix}.$
Of course, we also have
$\begin{pmatrix}\int_{\mathbb{R}^{d}}\psi_{1}\textnormal{d}H_{0}\\ \vdots\\ \int_{\mathbb{R}^{d}}\psi_{k}\textnormal{d}H_{0}\end{pmatrix}=\begin{pmatrix}\int_{\mathbb{R}^{d}}\psi_{1}\textnormal{d}G\\ \vdots\\ \int_{\mathbb{R}^{d}}\psi_{k}\textnormal{d}G\end{pmatrix}$
by construction. Since $A$ being strictly convex implies that $A^{\ast}$ is continuously differentiable, we conclude $\alpha_{n}\to 0$ as $n\to\infty$. Also, the relative entropy of $H_{\alpha_{n}}$ from $H_{0}$ is just the $A$-Bregman divergence of $\alpha_{n}$ from 00, so combining $\alpha_{n}\to 0$ and Pinsker’s inequality shows that we have $\|H_{\alpha_{n}}-H_{0}\|_{\operatorname{TV}}\to 0$ as $n\to\infty$. Since $\{H_{\alpha}\}_{\alpha}$ have common compact support, this further implies $W_{2}(H_{\alpha_{n}},H)\to 0$. Now use $W_{2}(\bar{F}_{n},F)\to 0$ and the gluing lemma Villani [2009, p. 11] to construct $\pi_{n}\in\Gamma(\bar{F}_{n};H_{\alpha_{n}})$ for all $n\in\mathbb{N}$ such that we have $W_{2}(\pi_{n},\pi)\to 0$. By construction, this means we have $\pi_{n}\in\Gamma_{K}(\bar{F}_{n};\hat{G}_{n},\Psi)$, so the proof is complete. ∎
Proof of Theorem C.5.
We claim (as stated in the discussion in Section 4.3) that we have $W_{2}(\hat{H}_{n},H)\to 0$ in probability as $n\to\infty$. If this is true, then we may apply the exact same argument from the proof of Theorem 4.4 so we are done.
In fact, we will prove a slightly stronger claim. For convenience, write $\hat{\pi}_{n}$ for $\hat{\pi}_{\mathcal{GCB}}$, in order to emphasize the dependence on $n\in\mathbb{N}$. Then, we claim that we have $W_{2}(\hat{\pi}_{n},\pi_{\mathcal{GCB}})\to 0$ in probability as $n\to\infty$. Since assumptions (ZC), (BC), (C), and (QG) imply (Theorem C.3) that $\pi_{\mathcal{GCB}}$ is the unique solution to (3.4), it suffices to show that $\{\hat{\pi}_{n}\}_{n\in\mathbb{N}}$ is $W_{2}$-precompact and that every subsequential $W_{2}$-limit is a solution to (3.4). To see pre-compactness, simply note that the marginal distributions of $\{\hat{\pi}_{n}\}_{n\in\mathbb{N}}$ are $\{\bar{F}_{n}\}_{n\in\mathbb{N}}$ and $\{\hat{H}_{n}\}_{n\in\mathbb{N}}$; the former converges in $W_{2}$ to $F$ hence it is $W_{2}$-pre-compact, and the latter is $W_{2}$-pre-compact since these probability measures have uniformly bounded support.
Next, suppose that $\{n_{k}\}_{k\in\mathbb{N}}$ and $\pi_{\infty}\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ have $W_{2}(\hat{\pi}_{\mathcal{DCB},n_{k}},\pi_{\infty})\to 0$, and let us show that $\pi_{\infty}$ is optimal for (3.4). To see that $\pi_{\infty}$ is feasible for (3.4), take arbitrary $1\leq\ell\leq k$. Then, use the fact that $\{\hat{H}_{n}\}_{n\in\mathbb{N}}$ have uniformly bounded support, the fact that $\hat{H}_{n}\in\Gamma(\bar{F}_{n};\hat{G}_{n},\Psi)$ for all $n\in\mathbb{N}$, and assumption (CC) to get:
$\int_{\mathbb{R}^{d}\times K}\psi_{\ell}(\eta)\textnormal{d}\pi_{\infty}(z,\eta)$ $=\lim_{n\to\infty}\int_{\mathbb{R}^{d}\times K}\psi_{\ell}(\eta)\textnormal{d}\hat{\pi}_{n}(z,\eta)$
$=\lim_{n\to\infty}\int_{K}\psi_{\ell}(\eta)\textnormal{d}\hat{G}_{n}(\eta)=\int_{K}\psi_{\ell}(\eta)\textnormal{d}G(\eta).$
Also, the first coordinate of $\pi_{\infty}$ is the $W_{2}$-limit of the first coordinates $\bar{F}_{n}$ of $\hat{\pi}_{n}$, which is $F$. Thus, $\pi_{\infty}$ is feasible.
Next, we develop a useful estimate. We claim that any $\{\pi_{n}\}_{n\in\mathbb{N}}$ with $\pi_{n}\in\Gamma(\bar{F}_{n};\hat{H}_{n})$ must satisfy
$\left|\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi_{n}-\int_{\mathbb{R}^{d}\times K}\hat{c}_{n}\textnormal{d}\pi_{n}\right|\to 0.$ (G.17)
as $n\to\infty$. To show this, we calculate:
$\left|\int_{\mathbb{R}^{d}\times K}c_{G}\textnormal{d}\pi_{n}-\int_{\mathbb{R}^{d}\times K}\hat{c}_{n}\textnormal{d}\pi_{n}\right|$
$=\left|\int_{\mathbb{R}^{d}\times K}\left(\|\delta_{\mathcal{B}}(z)-\eta\|^{2}-\|\hat{\delta}_{\mathcal{B}}(z)-\eta\|^{2}\right)\textnormal{d}\pi_{n}(z,\eta)\right|$
$=\left|\int_{\mathbb{R}^{d}\times K}\left(\|\delta_{\mathcal{B}}(z)\|^{2}-\|\hat{\delta}_{\mathcal{B}}(z)\|^{2}-2\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\eta\right\rangle\right)\textnormal{d}\pi_{n}(z,\eta)\right|$
$\leq\left|\int_{\mathbb{R}^{d}\times K}\left(\|\delta_{\mathcal{B}}(z)\|^{2}-\|\hat{\delta}_{\mathcal{B}}(z)\|^{2}\right)\textnormal{d}\pi_{n}(z,\eta)\right|+2\left|\int_{\mathbb{R}^{d}\times K}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\eta\right\rangle\textnormal{d}\pi_{n}(z,\eta)\right|$
$\leq\int_{\mathbb{R}^{d}\times K}\left|\|\delta_{\mathcal{B}}(z)\|^{2}-\|\hat{\delta}_{\mathcal{B}}(z)\|^{2}\right|\textnormal{d}\pi_{n}(z,\eta)+2\int_{\mathbb{R}^{d}\times K}\left|\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\eta\right\rangle\right|\textnormal{d}\pi_{n}(z,\eta),$
and we separately further the bound on each of these terms. For the first term, we notice that the integral only depends on the marginal distribution of $z$ under $\pi_{n}$, which equals $\bar{F}_{n}$ by construction. Also, we have the elementary inequality
$\left|\|x\|^{2}-\|y\|^{2}\right|\leq\|x-y\|^{2}+2\|x\|\cdot\|x-y\|$
for all $x,y\in\mathbb{R}^{m}$. So, taking $x=\delta_{\mathcal{B}}(Z_{i})$ and $y=\hat{\delta}_{\mathcal{B}}(Z_{i})$, summing, and applying Jensen’s inequality yields
$\int_{\mathbb{R}^{d}\times K}\left|\|\delta_{\mathcal{B}}(z)\|^{2}-\|\hat{\delta}_{\mathcal{B}}(z)\|^{2}\right|\textnormal{d}\pi_{n}(z,\eta)$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left|\|\delta_{\mathcal{B}}(Z_{i})\|^{2}-\|\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}\right|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left(\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}+2\|\delta_{\mathcal{B}}(Z_{i})\|\cdot\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}+\frac{2}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})\|\cdot\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}+2\cdot\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\|^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\delta_{\mathcal{B}}(Z_{i})\|^{2}}.$
The the right side above vanishes as $n\to\infty$ because of the assumption (EBQ), assumption (QG), and the law of large numbers. For the second term above, use Cauchy-Schwarz twice to get
$\int_{\mathbb{R}^{d}\times K}\left|\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\eta\right\rangle\right|\textnormal{d}\pi_{n}(z,\eta)$
$\leq\int_{\mathbb{R}^{d}\times K}\left\|\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z)\right\|\cdot\|\eta\|\textnormal{d}\pi_{n}(z,\eta)$
$\leq\sqrt{\int_{\mathbb{R}^{d}\times K}\left\|\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z)\right\|\textnormal{d}\pi_{n}(z,\eta)}\sqrt{\int_{\mathbb{R}^{d}\times K}\|\eta\|^{2}\textnormal{d}\pi_{n}(z,\eta)}$
$=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left\|\delta_{\mathcal{B}}(Z_{i})-\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|^{2}}\sqrt{\int_{K}\|\eta\|^{2}\textnormal{d}\hat{H}_{n}(\eta)}.$
As before, the first term vanishes because of (EBQ) and the second term is bounded because of (CS). Thus, we have shown (G.17).
Lastly, we show that $\pi_{\infty}$ is in fact optimal for (3.4). To do this, take an arbitrary $\pi\in\Gamma(F;H)$ and use Lemma G.2 to get some $\{\pi_{k}\}_{k\in\mathbb{N}}$ with $\pi_{k}\in\Gamma(\bar{F}_{n_{k}};\hat{H}_{n_{k}})$ and $W_{2}(\pi_{k},\pi)\to 0$ as $k\to\infty$. Then use the Portmanteau lemma, the estimate (G.17), the optimality of $\{\hat{\pi}_{n_{k}}\}_{k\in\mathbb{N}}$, and the estimate (G.17) again, to get:
$\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi^{\ast}$ $\leq\liminf_{k\to\infty}\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\hat{\pi}_{n_{k}}$
$\leq\liminf_{k\to\infty}\int_{\mathbb{R}^{d}\times K}\hat{c}_{n_{k}}\,\textnormal{d}\hat{\pi}_{n_{k}}$
$\leq\liminf_{k\to\infty}\int_{\mathbb{R}^{d}\times K}\hat{c}_{n_{k}}\,\textnormal{d}\pi_{k}$
$\leq\liminf_{k\to\infty}\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi_{k}$
In fact, we have
$\lim_{k\to\infty}\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi_{k}=\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi$
by $W_{2}$ convergence, since $c_{G}$ is uniformly integrable under $\{\pi_{k}\}_{k\in\mathbb{N}}$. Thus $\pi_{\infty}$ is optimal, hence $\pi_{\infty}=\pi_{\mathcal{GCB}}$. This completes the proof. ∎